### v1

# Collective order boundedness of sets of operators between ordered vector spaces

Abstract
It is proved that: each collectively order continuous set of operators from an Archimedean ordered vector space to an ordered vector space is collectively order bounded; each collectively order to norm bounded set of operators from an ordered Banach space with a closed generating cone to a normed space is norm bounded; collectively r-continuous sets agree with collectively order bounded ones when the space in the range has a generating cone. Several applications to operator semigroups are given.
Keywords: ordered vector space, collectively qualified set of operators, operator semigroup
MSC2020: 46A40, 46B42, 47B60, 47D03
1 Introduction
Collectively compact sets of operators between normed spaces were introduced by P. M. Anselone and T. W. Palmer [5]. Recently, collectively qualified sets of operators on Banach latticed were investigate in [4, 7, 8, 9]. The present paper generalizes results of [7] to the ordered vector space setting. The results of the paper are new even in the single operator case. For example, they imply that: 1) each order continuous operator between Archimedean ordered vector spaces is order bounded; 2) each order to norm continuous operator from an ordered Banach space with a closed generating cone to a normed space is continuous.
We shall abbreviate a normed space (ordered vector space, ordered normed space, ordered Banach space, vector lattice, normed lattice, Banach lattice) by NS (OVS, ONS, OBS, VL, NL, BL). Throughout the text, vector spaces are real, operators are linear, symbol ${\cal L}(X,Y)$ (${\cal L}_{+}(X,Y)$, ${\cal L}_{ob}(X,Y)$, ${\cal L}_{oc}(X,Y)$, ${\cal L}_{onb}(X,Y)$) stands for the space of all (positive, order bounded, order continuous, order to norm bounded) operators from $X$ to $Y$. A symbol $x_{\alpha}\downarrow 0$ stands for a decreasing net in an OVS such that $\inf\limits_{\alpha}x_{\alpha}=0$.
Quite often, working with collective properties of a set of operators in OVSs involves order convergent nets. Let $X$ be an OVS. A net $(x_{\alpha})$ in $X$ is
  1. -
    order convergent to $x\in X$ (briefly, o-convergent to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{o}}x$) whenever there exists a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$ there is an $\alpha_{\beta}$ satisfying $-g_{\beta}\leq x_{\alpha}-x\leq g_{\beta}$ for all $\alpha\geq\alpha_{\beta}$.
  2. -
    relative uniform convergent to $x\in X$ (briefly, r-convergent to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{r}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices satisfying $-\frac{1}{n}u\leq x_{\alpha}-x\leq\frac{1}{n}u$ for all $\alpha\geq\alpha_{n}$.
We shall use the following OVS version of [7, Definition 2.1]. Let $X$ be an OVS, $B$ a set, $A$ a directed set, and =${\cal B}=\{(x^{b}_{\alpha})_{\alpha\in A}\}_{b\in B}$ a set of nets in $X$ indexed by $A$. We say that ${\cal B}$
  1. -
    collective o-converges to 00 (briefly, ${\cal B}\operatorname{\xrightarrow[]{c-o}}0$) if there exists a net $g_{\beta}\downarrow 0$ such that, for each $\beta$, there is an $\alpha_{\beta}$ satisfying ±$\pm x^{b}_{\alpha}\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$ and $b\in B$.
  2. -
    collective r-converges to 00 (briefly, ${\cal B}\operatorname{\xrightarrow[]{c-r}}0$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices with ±$\pm x^{b}_{\alpha}\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$ and $b\in B$.
[…]
The following definition extends [7, Definition 2.1] to OVSs.
Definition 1.1.
Let ${\cal T}\subseteq{\cal L}(X,Y)$.
When $X$ and $Y$ are OVSs,
  1. -
    ${\cal T}$ is collectively order bounded (briefly, ${\cal T}\in{\bf L}_{ob}(X,Y)$) if, for every order interval $[a,b]$ in $X$ there exists an order interval $[u,v]$ in $Y$ such that ${\cal T}[a,b]:=\bigcup\limits_{T\in\mathcal{T}}{T}[a,b]\subseteq[u,v]$.
  2. -
    ${\cal T}$ is collectively o-continuous (briefly, ${\cal T}\in\text{\bf L}_{oc}(X,Y)$) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c-o}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$;
  3. -
    ${\cal T}$ is collectively r-continuous (briefly, ${\cal T}\in\text{\bf L}_{rc}(X,Y)$) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c-r}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{r}}0$;
When $X$ is an OVS and $Y$ is an ONS,
  1. -
    ${\cal T}$ is collectively almost order bounded (briefly, ${\cal T}\in\text{\bf L}_{aob}(X,Y)$) if, for every $[a,b]\subseteq X$ and $\varepsilon>0$, there exists $[u,v]\subseteq Y$ such that ${\cal T}[a,b]\subseteq[u,v]+\varepsilon B_{Y}$.
When $X$ is a normed space and $Y$ is an ONS,
  1. -
    ${\cal T}$ is collectively semicompact (briefly, ${\cal T}\in\text{\bf L}_{sc}(X,Y)$) if, for every $\varepsilon>0$, there exists $[u,v]\subseteq Y$ such that ${\cal T}B_{X}\subseteq[u,v]+\varepsilon B_{Y}$.
When $X$ is an OVS and $Y$ is a normed space,
  1. -
    ${\cal T}$ is collectively order to norm bounded (briefly, ${\cal T}\in\text{\bf L}_{onb}(X,Y)$) if the set ${\cal T}[a,b]$ is norm bounded for every $[a,b]\subseteq X$.
The present paper is organized as follows. Section 2 begins with Theorem 2.1 saying that every collectively order continuous set of operators from an OVS to an Archimedean OVS is collectively order bounded. Then, we prove Theorem 2.2 that describes relationship between collectively r-continuous and collectively order bounded sets. Finally, we prove Theorem 2.3 saying that under rather mild conditions, every collectively order to norm bounded set of operators is norm bounded. Section 3 is devoted to applications of obtained results to one-parameter operator semigroups on OVSs.
For the terminology and notations not explained in the text, we refer to [2, 3, 6, 10, 11].
2 Main results
We are going to prove the following results which extend the results obtained recently in [7] in the vector lattice setting.
  1. -
    Every collectively order continuous set of operators from an Archimedean OVS to an OVS is collectively order bounded.
  1. -
    Every collectively order to norm bounded set of operators from an OBS with a closed generating cone to a NS is norm bounded
  1. -
    Collectively r-continuous sets coincides with collectively order bounded sets when the OVS in the range has a generating cone.
2.1
We begin with an extension of the first part of [7, Theorem 2.2].
Theorem 2.1.
Let $X$ and $Y$ be OVSs, with an Archimedean $X$. Then $\text{\bf L}_{oc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$.
Proof.
Let ${\cal T}\in\text{\bf L}_{oc}(X,Y)$. It is enough to prove order boundedness of ${\cal T}[0,b]$ for every $b\in X_{+}$. Take $[0,b]\subseteq X$. Following [1, Theorem 2.1], consider the lexicographically directed set =×${\cal I}=\mathbb{N}\times[0,b]$: $(m,z)\geq(n,z^{\prime})$ if $m>n$ or $m=n$ and $z\geq z^{\prime}$, and let $x_{(k,z)}=\frac{1}{k}z$. Since $X$ is Archimedean, $\frac{1}{n}b\downarrow 0$. It follows from $0\leq x_{(m,z)}=\frac{1}{m}z\leq\frac{1}{n}b$ for $(m,z)\geq(n,b)$ that $x_{(m,z)}\operatorname{\xrightarrow[]{o}}0$.
Since ${\cal T}\in\text{\bf L}_{oc}(X,Y)$, there exists $g_{\beta}\downarrow 0$ in $Y$ such that, for every $\beta$ there exists $(m_{\beta},y_{\beta})$ satisfying ±$\pm Tx_{(m,y)}\leq g_{\beta}$ for all $T\in{\cal T}$ and $(m,y)\geq(m_{\beta},y_{\beta})$. Pick any $g_{\beta_{0}}$. Since $(m_{\beta_{0}}+1,y)\geq(m_{\beta_{0}},y_{\beta_{0}})$,
±$\pm T\Big{(}\frac{y}{m_{\beta_{0}}+1}\Big{)}=\pm Tx_{(m_{\beta_{0}}+1,y)}\leq g_{\beta_{0}}\ \ \ (y\in[0,b],T\in{\cal T}).$
Therefore, ${\cal T}[0,b]\subseteq[-(m_{\beta_{0}}+1)g_{\beta_{0}},(m_{\beta_{0}}+1)g_{\beta_{0}}]$. ∎
In particular, Theorem 2.1 generalizes the Abramovich–Sirotkin theorem [1, Theorem 2.1] in the following way.
Corollary 2.1.
Each order continuous operator from an Archimedean OVS to an OVS is order bounded.
The Archimedean assumption is essential in Corollary 2.1.
Example 2.1.
Let $X$ be the lexicographically ordered real plane $\mathbb{R}^{2}$, $Y$ the pointwise ordered $\mathbb{R}^{2}$, and $T:X\to Y$ the identity operator. Both $X$ and $Y$ are VLs. The operator $T$ is not order bounded since the order interval $[(0,0),(2,0)]$ of $X$ contains line $x=1$ that is not order bounded in $Y$. However, $T$ is order continuous. To see this, it is enough to show that $(x_{\alpha},y_{\alpha})\downarrow 0$ in $X$ implies $T((x_{\alpha},y_{\alpha}))=(x_{\alpha},y_{\alpha})\operatorname{\xrightarrow[]{o}}0$ in $Y$. Suppose $(x_{\alpha},y_{\alpha})\downarrow 0$ in $X$ and let $\varepsilon>0$. There exists $\alpha_{0}$ with $x_{\alpha}=0$ for $\alpha\geq\alpha_{0}$. Now, take $\alpha_{1}\geq\alpha_{0}$ such that $0\leq\max\{x_{\alpha},y_{\alpha}\}=y_{\alpha}\leq\varepsilon$ for $\alpha\geq\alpha_{1}$. It follows that $T(x_{\alpha},y_{\alpha})\operatorname{\xrightarrow[]{o}}0$ in $Y$.
2.2
Next, we extend the last part of [7, Theorem 2.2] to OVSs.
Theorem 2.2.
Let $X$ and $Y$ be OVSs then $\text{\bf L}_{rc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$. If additionally $Y=Y_{+}-Y_{+}$ then $\text{\bf L}_{rc}(X,Y)=\text{\bf L}_{ob}(X,Y)$.
Proof.
Let ${\cal T}\in\text{\bf L}_{rc}(X,Y)$. In order to show ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ it suffices to prove that ${\cal T}[0,b]$ is order bounded in $Y$ for each $b\in X_{+}$. Take $[0,b]\subseteq X$. Consider =×${\cal I}=\mathbb{N}\times[0,b]$ and the net $x_{(k,z)}=\frac{1}{k}z$ like in the proof of Theorem 2.1. If $(m,z)\geq(n,b)$ in ${\cal I}$ then either $(m,z)=(n,b)$ or $m>n$, and hence $0\leq x_{(m,z)}=\frac{1}{m}z\leq\frac{1}{n}b$. Thus, $x_{(m,z)}\operatorname{\xrightarrow[]{r}}0$ in $X$.
Since ${\cal T}\in\text{\bf L}_{rc}(X,Y)$, there exist $u\in Y_{+}$ and an increasing sequence $(k_{n},z_{n})$ in ${\cal I}$ satisfying ±$\pm T\big{(}\frac{z}{k}\big{)}=\pm Tx_{(k,z)}\leq\frac{1}{n}u$ for all $T\in{\cal T}$ and $(k,z)\geq(k_{n},z_{n})$. In particular, ±$\pm T\big{(}\frac{z}{k_{1}+1}\big{)}=\pm Tx_{(k_{1}+1,z)}\leq u$ for all $T\in{\cal T}$ and $z\in[0,b]$. So, ${\cal T}[0,b]\subseteq[-(k_{1}+1)u,(k_{1}+1)u]$. It follows ${\cal T}\in\text{\bf L}_{ob}(X,Y)$.
Now, suppose $Y=Y_{+}-Y_{+}$. Let ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ and $x_{\alpha}\operatorname{\xrightarrow[]{r}}0$. Then, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices such that $-u\leq nx_{\alpha}\leq u$ (and hence $0\leq u+nx_{\alpha}\leq 2u$) for all $\alpha\geq\alpha_{n}$. Since ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ and $Y=Y_{+}-Y_{+}$, there exist $w_{1},w_{2}\in Y_{+}$ satisfying $Tu\in[-w_{1},w_{1}]$ and $T(u+nx_{\alpha})\in[-w_{2},w_{2}]$ for all $T\in{\cal T}$ and $\alpha\geq\alpha_{n}$. Then $nTx_{\alpha}\in[-(w_{1}+w_{2}),(w_{1}+w_{2})]$ (and hence ±$\pm Tx_{\alpha}\leq\frac{w_{1}+w_{2}}{n}$) for $T\in{\cal T}$ and $\alpha\geq\alpha_{n}$. It follows ${\cal T}\in\text{\bf L}_{rc}(X,Y)$. ∎
We do not know whether the assumption $Y=Y_{+}-Y_{+}$ in Theorem 2.2 is essential for $\text{\bf L}_{rc}(X,Y)=\text{\bf L}_{ob}(X,Y)$.
Corollary 2.2.
Order bounded operators from an OVS to an OVS with generating positive cone agree with r-continuous operators.
2.3
Recall that order intervals in an OBS $Y$ with a closed cone are norm bounded iff the cone is normal (cf., [3, Theorem 2.40]). We include the following elementary collective extension of this fact .
Proposition 2.1.
Let $X$ be an OVS and $Y$ be an OBS such that $Y_{+}$ is closed, generating, and normal. Then $\text{\bf L}_{aob}(X,Y)\subseteq\text{\bf L}_{onb}(X,Y)$.
Proof.
Let ${\cal T}\in\text{\bf L}_{aob}(X,Y)$ and take $[a,b]\subseteq X$. Then $\bigcup_{T\in\mathcal{T}}{T}[a,b]\subseteq[-y_{1},y_{1}]+B_{Y}$ for some $y_{1}\in Y_{+}$. Since $Y_{+}$ is normal, $[-y_{1},y_{1}]+B_{Y}$ is norm bounded by [3, Theorem 2.40]. Therefore, ${\cal T}\in\text{\bf L}_{onb}(X,Y)$. ∎
The normality of $Y_{+}$ is essential in Proposition 2.1. Indeed, the identity operator in the OBS $C^{1}[0,1]$ of continuously differentiable functions on $[0,1]$ is order bounded yet not order to norm bounded.
The following result generalizes [7, Theorem 2.1] to OBSs.
Theorem 2.3.
Let $X$ be an ordered Banach space with a closed generating cone and $Y$ be a normed space. Then each ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ is norm bounded.
Proof.
Let ${\cal T}\in\text{\bf L}_{onb}(X,Y)$. Suppose in contrary ${\cal T}B_{X}$ is not norm bounded. The M. G. Krein – Smulian theorem (cf., [3, Theorem 2.37]) implies $\alpha B_{X}\subseteq B_{X}\cap X_{+}-B_{X}\cap X_{+}$ for some $\alpha>0$, and hence ${\cal T}(B_{X}\cap X_{+})$ is not norm bounded. Pick sequences $(x_{n})$ in $B_{X}\cap X_{+}$ and $(T_{n})$ in ${\cal T}$ with $\|T_{n}x_{n}\|\geq n^{3}$ for all $n$, and set $x:=\|\cdot\|-$$\sum\limits_{n=1}^{\infty}n^{-2}x_{n}\in X_{+}$. Since ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ then ${\cal T}[0,x]\subseteq MB_{Y}$ for some $M<\infty$. It follows from $n^{-2}x_{n}\in[0,x]$ that $\|T_{n}(n^{-2}x_{n})\|\leq M$ for every $n$. This is absurd, because $\|T_{n}(n^{-2}x_{n})\|\geq n$ for all $n$. The proof is complete. ∎
The condition $X_{+}-X_{+}=X$ is essential in Theorem 2.3.
Example 2.2.
Take any infinite dimensional Banach space $X$ with the trivial cone $\{0\}$, consider arbitrary discontinuous operator $T$ on $X$, and observe that $T$ is order to norm bounded simply because order intervals in $X$ are singletons.
3 Applications to commutative operator semigroups
We shall use the notation $(T_{s})_{s\in J}$ for a one-parameter operator semigroup on a vector space $X$ (with $J=\mathbb{R}_{+}$ in the continuous-parameter case and $J=\mathbb{Z}_{+}$ in the discrete-parameter case). More generally, any commutative operator semigroup $\mathcal{G}$ on $X$ can be written as $\mathcal{G}=(T_{S})_{S\in J}$, where $T_{S}=S$ and $J=\mathcal{G}$ is directed by a partial order $RS\succeq S$ for all $R,S\in\mathcal{G}$. The partial order $\succeq$ is extendable to the commutative operator semigroup ${\text{\rm co}}(\mathcal{G})$ via letting $RS\succeq S$ for all $R,S\in{\text{\rm co}}(\mathcal{G})$ (cf., [10, p.75]).
Proposition 3.1.
Let $\mathcal{G}$ be a collectively order to norm bounded commutative operator semigroup on an OBS $X$ containing a (weakly) compact operator $S_{0}$. Then the semigroup $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively (weakly) compact. In particular, $\{S\}_{S\succeq S_{0}}$ is collectively (weakly) compact.
Proof.
By Theorem 2.3, $\|S\|\leq M<\infty$ for $S\in\mathcal{G}$, and hence for all $S\in{\text{\rm co}}(\mathcal{G})$. Thus,
$\bigcup_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}SB_{X}=S_{0}\Big{(}\bigcup_{S\in{\text{\rm co}}(\mathcal{G})}SB_{X}\Big{)}\subseteq S_{0}(MB_{X})=MS_{0}B_{X}.$
Since $S_{0}B_{X}$ is relatively (weakly) compact, $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively (weakly) compact. ∎
Combining Propositions 3.1 and 2.1 we obtain.
Corollary 3.1.
Let $\mathcal{G}$ be a collectively almost order bounded commutative operator semigroup on an OBS $X$ with a closed, generating, and normal cone. If $\mathcal{G}$ contains a (weakly) compact operator $S_{0}$ then $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively (weakly) compact. In particular, $\{S\}_{S\succeq S_{0}}$ is collectively (weakly) compact.
Next, we consider commutative operator semigroups on AL-spaces.
Proposition 3.2.
Let $\mathcal{G}$ be a collectively almost order bounded commutative operator semigroup on on an AL-space $E$ containing a weakly compact operator $S_{0}$. Then the semigroup $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively semicompact. In particular, $\{S\}_{S\succeq S_{0}}$ is collectively semicompact.
Proof.
By Theorem 2.3, $\|S\|\leq M<\infty$ for $S\in\mathcal{G}$. Then $\|S\|\leq M<\infty$ for all $S\in{\text{\rm co}}(\mathcal{G})$. We conclude $\bigcup_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}SB_{E}\subseteq MS_{0}B_{E}$. Since the set $S_{0}B_{E}$ is relatively weakly compact, it is almost order bounded by the Dunford-Pettis theorem (cf., [11][Theorem 2.5.4]). Thus, $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively semicompact. ∎
Combining Propositions 3.2 and 2.1 we obtain.
Corollary 3.2.
Let $\mathcal{G}$ be a collectively almost order bounded commutative operator semigroup on an AL-space $E$ containing a weakly compact operator $S_{0}$. Then the semigroup $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively semicompact. In particular, $\{S\}_{S\succeq S_{0}}$ is collectively semicompact.
Finally, we include the following convergence theorem.
Theorem 3.1.
Let $\mathcal{G}$ be a collectively almost order bounded commutative operator semigroup on an order continuous BL $E$. Then the net $(S)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ converges strongly to a projection.
Proof.
By Theorem 2.3, $\|S\|\leq M<\infty$ for $S\in\mathcal{G}$, and hence for all $S\in\overline{{\text{\rm co}}}(\mathcal{G})$. Let $x\in E_{+}$. By the assumption, the set $\bigcup_{S\in\mathcal{G}}S[-x,x]$ is almost order bounded. Since $E$ is order continuous, $\bigcup_{S\in\mathcal{G}}S[-x,x]$ is relatively weakly compact due to the Grothendieck lemma (cf., [2, Theorem 3.44]).
Since $(Sx)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ lies in the weakly compact set $\overline{\text{\rm co}}(\bigcup_{S\in\mathcal{G}}S[-x,x])$, it has a weakly cluster point, say $y$. Then $\lim\limits_{S\to\infty}\|Sx-y\|=0$ by the Eberlein theorem (cf., [10, Theorem 1.5]). Since $x\in E_{+}$ is arbitrary, we conclude that $(S)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ converges strongly to a projection. ∎
Corollary 3.3.
Let $(T_{s})_{s\in J}$ be either a $C_{0}$-semigroup or a discrete operator semigroup on an order continuous BL $E$. If $(T_{s})_{s\in J}$ is collectively almost order bounded then $(T_{s})_{s\in J}$ is mean ergodic.

### v2

@@
 Abstract
-It is proved that: each collectively order continuous set of operators from an Archimedean ordered vector space to an ordered vector space is collectively order bounded; each collectively order to norm bounded set of operators from an ordered Banach space with a closed generating cone to a normed space is norm bounded; collectively r-continuous sets agree with collectively order bounded ones when the space in the range has a generating cone.
-Several applications to operator semigroups are given.
-Keywords: ordered vector space, collectively qualified set of operators, operator semigroup
+It is proved that: each collectively order continuous set of operators between Archimedean ordered vector spaces with generating cones is collectively order bounded; and each collectively order to norm bounded set of operators from an ordered Banach space with a closed generating cone to a normed space is norm bounded.
+Several applications to commutative operator semigroups on ordered vector spaces are given.
+Keywords: ordered vector space, collectively qualified set of operators, commutative operator semigroup
 MSC2020: 46A40, 46B42, 47B60, 47D03
@@
 Recently, collectively qualified sets of operators on Banach latticed were investigate in [4, 7, 8, 9].
-The present paper generalizes results of [7] to the ordered vector space setting.
+The present paper generalizes results of [1, 7, 15] to the ordered vector space setting.
 The results of the paper are new even in the single operator case.
@@
   2. -
-    relative uniform convergent to $x\in X$ (briefly, r-convergent to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{r}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices satisfying $-\frac{1}{n}u\leq x_{\alpha}-x\leq\frac{1}{n}u$ for all $\alpha\geq\alpha_{n}$.
+    relative uniform convergent to $x\in X$ (briefly, ru-convergent to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{r}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices satisfying $-\frac{1}{n}u\leq x_{\alpha}-x\leq\frac{1}{n}u$ for all $\alpha\geq\alpha_{n}$.
 We shall use the following OVS version of [7, Definition 2.1].
@@
   2. -
-    collective r-converges to 00 (briefly, ${\cal B}\operatorname{\xrightarrow[]{c-r}}0$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices with ±$\pm x^{b}_{\alpha}\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$ and $b\in B$.
+    collective ru-converges to 00 (briefly, ${\cal B}\operatorname{\xrightarrow[]{c-r}}0$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices with ±$\pm x^{b}_{\alpha}\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$ and $b\in B$.
 […]
@@
   3. -
-    ${\cal T}$ is collectively r-continuous (briefly, ${\cal T}\in\text{\bf L}_{rc}(X,Y)$) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c-r}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{r}}0$;
+    ${\cal T}$ is collectively ru-continuous (briefly, ${\cal T}\in\text{\bf L}_{rc}(X,Y)$) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c-r}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{r}}0$;
 When $X$ is an OVS and $Y$ is an ONS,
@@
 Then, we prove Theorem 2.2 that describes relationship between collectively r-continuous and collectively order bounded sets.
-Finally, we prove Theorem 2.3 saying that under rather mild conditions, every collectively order to norm bounded set of operators is norm bounded.
-Section 3 is devoted to applications of obtained results to one-parameter operator semigroups on OVSs.
+Finally, we prove Theorem 2.3 saying that under rather mild assumptions, every collectively order to norm bounded set of operators is norm bounded.
+It follows, in particular that under these assumptions each operator which maps order interval into balls is bounded.
+Section 3 is devoted to applications of obtained results to commutative operator semigroups on OVSs.
 For the terminology and notations not explained in the text, we refer to [2, 3, 6, 13, 14].
 2 Main results
-We are going to prove the following results which extend the results obtained recently in [7] in the vector lattice setting.
+We are going to prove the following results which generalize the results obtained recently in [7, 15] in the vector lattice setting to the OVS setting.
   1. -
-    Every collectively order continuous set of operators from an Archimedean OVS to an OVS is collectively order bounded.
+    Every collectively order continuous set of operators from an Archimedean OVS with a generating cone to an OVS with a generating cone is collectively order bounded.
   1. -
@@
 2.1
-We begin with an extension of the first part of [7, Theorem 2.2].
+We begin with generalization of the first part of [7, Theorem 2.2]. The proof of the next theorem is based on ideas of the proof of the Abramovich–Sirotkin theorem in [1].
 Theorem 2.1.
-Let $X$ and $Y$ be OVSs, with an Archimedean $X$. Then $\text{\bf L}_{oc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$.
+Let $X$ and $Y$ be OVSs with generating cones and suppose $X$ is Archimedean. Then $\text{\bf L}_{oc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$.
 Proof.
@@
 Therefore, ${\cal T}[0,b]\subseteq[-(m_{\beta_{0}}+1)g_{\beta_{0}},(m_{\beta_{0}}+1)g_{\beta_{0}}]$. ∎
-In particular, Theorem 2.1 generalizes the Abramovich–Sirotkin theorem [1, Theorem 2.1] in the following way.
+In particular, Theorem 2.1 generalizes the Abramovich–Sirotkin theorem [1, Theorem 2.1] as follows.
 Corollary 2.1.
-Each order continuous operator from an Archimedean OVS to an OVS is order bounded.
-The Archimedean assumption is essential in Corollary 2.1.
+Each order continuous operator from an Archimedean OVS $X$ with generating cone to an OVS $Y$ with generating cone is order bounded.
+The following example shows that the Archimedean assumption, which has being missed in [1, Theorem 2.1], is essential in Corollary 2.1 even in the VL setting.
 Example 2.1.
@@
 2.2
-Next, we extend the last part of [7, Theorem 2.2] to OVSs.
+Next, we generalize [15, Theorem 5.1] and [7, Theorem 2.2] to OVSs.
 Theorem 2.2.
-Let $X$ and $Y$ be OVSs then $\text{\bf L}_{rc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$. If additionally $Y=Y_{+}-Y_{+}$ then $\text{\bf L}_{rc}(X,Y)=\text{\bf L}_{ob}(X,Y)$.
+Let $X$ and $Y$ be OVSs with generating cones then $\text{\bf L}_{rc}(X,Y)=\text{\bf L}_{ob}(X,Y)$.
 Proof.
@@
 ∎
-We do not know whether the assumption $Y=Y_{+}-Y_{+}$ in Theorem 2.2 is essential for $\text{\bf L}_{rc}(X,Y)=\text{\bf L}_{ob}(X,Y)$.
 Corollary 2.2.
-Order bounded operators from an OVS to an OVS with generating positive cone agree with r-continuous operators.
+Order bounded operators between OVSs with generating cones agree with r-continuous operators.
 2.3
@@
 Proposition 2.1.
-Let $X$ be an OVS and $Y$ be an OBS such that $Y_{+}$ is closed, generating, and normal. Then $\text{\bf L}_{aob}(X,Y)\subseteq\text{\bf L}_{onb}(X,Y)$.
+Let $X$ be an OVS and $Y$ be an OBS with closed generating normal cone. Then $\text{\bf L}_{aob}(X,Y)\subseteq\text{\bf L}_{onb}(X,Y)$.
 Proof.
@@
 Theorem 2.3.
-Let $X$ be an ordered Banach space with a closed generating cone and $Y$ be a normed space. Then each ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ is norm bounded.
+Let $X$ be an OBS with a closed generating cone and $Y$ be a normed space. Then each ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ is norm bounded.
 Proof.
@@
 ∎
+It is worth noting that even in the case of a single operator, Theorem 2.3 generalizes the well known fact that every order bounded operator between BLs is bounded.
 The condition $X_{+}-X_{+}=X$ is essential in Theorem 2.3.
@@
 Take any infinite dimensional Banach space $X$ with the trivial cone $\{0\}$, consider arbitrary discontinuous operator $T$ on $X$, and observe that $T$ is order to norm bounded simply because order intervals in $X$ are singletons.
-3 Applications to commutative operator semigroups
+3 Some applications to operator semigroups
 We shall use the notation $(T_{s})_{s\in J}$ for a one-parameter operator semigroup on a vector space $X$ (with $J=\mathbb{R}_{+}$ in the continuous-parameter case and $J=\mathbb{Z}_{+}$ in the discrete-parameter case).
 More generally, any commutative operator semigroup $\mathcal{G}$ on $X$ can be written as $\mathcal{G}=(T_{S})_{S\in J}$, where $T_{S}=S$ and $J=\mathcal{G}$ is directed by a partial order $RS\succeq S$ for all $R,S\in\mathcal{G}$.
-The partial order $\succeq$ is extendable to the commutative operator semigroup ${\text{\rm co}}(\mathcal{G})$ via letting $RS\succeq S$ for all $R,S\in{\text{\rm co}}(\mathcal{G})$ (cf., [10, p.75]).
+The partial order $\succeq$ is extendable to the commutative operator semigroup ${\text{\rm co}}(\mathcal{G})$ via letting $RS\succeq S$ for all $R,S\in{\text{\rm co}}(\mathcal{G})$ (cf., [13, p.75]).
 Proposition 3.1.
-Let $\mathcal{G}$ be a collectively order to norm bounded commutative operator semigroup on an OBS $X$ containing a (weakly) compact operator $S_{0}$.
+Let $\mathcal{G}$ be a collectively order to norm bounded commutative operator semigroup on an OBS $X$ with a closed generating cone containing a (weakly) compact operator $S_{0}$.
 Then the semigroup $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively (weakly) compact.
@@
 Since $S_{0}B_{X}$ is relatively (weakly) compact, $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively (weakly) compact. ∎
-Combining Propositions 3.1 and 2.1 we obtain.
+Combining Propositions 3.1 and 2.1, and keeping in mind that every OBS with a generating normal cone is Archimedean, we obtain.
 Corollary 3.1.
@@
 Since $(Sx)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ lies in the weakly compact set $\overline{\text{\rm co}}(\bigcup_{S\in\mathcal{G}}S[-x,x])$, it has a weakly cluster point, say $y$.
-Then $\lim\limits_{S\to\infty}\|Sx-y\|=0$ by the Eberlein theorem (cf., [10, Theorem 1.5]).
+Then $\lim\limits_{S\to\infty}\|Sx-y\|=0$ by the Eberlein theorem (cf., [13, Theorem 1.5]).
 Since $x\in E_{+}$ is arbitrary, we conclude that $(S)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ converges strongly to a projection.
@@
 Corollary 3.3.
-Let $(T_{s})_{s\in J}$ be either a $C_{0}$-semigroup or a discrete operator semigroup on an order continuous BL $E$.
+Let $(T_{s})_{s\in J}$ be either a $C_{0}$-semigroup or a one-parameter discrete operator semigroup on an order continuous BL $E$.
 If $(T_{s})_{s\in J}$ is collectively almost order bounded then $(T_{s})_{s\in J}$ is mean ergodic.
+Uniformly continuous semigroups is a starting point in the theory of $C_{0}$-semigroups. Here we discuss similar elementary notions concerning semigroups on OVSs.
+The next two lemmas follow directly from [7, Theorems 2.1 and 2.2] and from Theorems 2.1, 2.2, and 2.3.
+Lemma 3.1.
+Every collectively order to norm bounded set of operators on an (OBS) with a closed generating cone is norm bounded.
+Lemma 3.2.
+Let $X$ be and OVS with generating cone. Then $\text{\bf L}_{rc}(X)=\text{\bf L}_{ob}(X)$. If $X$ is additionally Archimedean then $\text{\bf L}_{oc}(X)\subseteq\text{\bf L}_{ob}(X)$.
+Next definition localizes several notions concerning semigrops.
+Definition 3.1.
+A semigroup $(T_{s})_{s\geq 0}$ on an OVS $X$ is
+  1. $a)$
+    locally order bounded if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{ob}(X)$ for every $M$.
+  2. $b)$
+    locally o-continuous if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{oc}(X)$ for every $M$.
+  3. $c)$
+    locally ru-continuous if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{rc}(X)$ for every $M$.
+A semigroup $(T_{s})_{s\geq 0}$ on a normed OVS $X$ is
+  1. $e)$
+    locally order to norm bounded if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{onb}(X)$
+    for every $M$.
+In several important cases, in each of the four items of Definition 3.1 the condition “for every $M$” can be replaced by “for some $M$”.
+Proposition 3.3.
+Let $(T_{s})_{s\geq 0}$ be a semigroup on an OVS $X$ such that $X_{+}-X_{+}=X$.
+  1. $i)$
+    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{ob}(X)$ for some $\varepsilon>0$, then $(T_{s})_{s\geq 0}$ is locally order bounded.
+  2. $ii)$
+    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{rc}(X)$ for some $\varepsilon>0$, then $(T_{s})_{s\geq 0}$ is locally ru-continuous.
+  3. $iii)$
+    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{oc}(X)$ for some $\varepsilon>0$, then $(T_{s})_{s\geq 0}$ is locally order continuous.
+  4. $iv)$
+    If $X$ is Banach space, $X_{+}$ is closed, and $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{onb}(X)$ for some $\varepsilon>0$, then $(T_{s})_{s\geq 0}$ is locally order to norm bounded.
+Proof.
+$i)$ Let $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{ob}(X)$ for some $\varepsilon>0$ and fix $M\geq 0$.
+Let $n\varepsilon\geq M$, and take arbitrary $a\leq b$ in $X$.
+It follows from $X_{+}-X_{+}=X$ that $[a,b]\subseteq[-a_{0},a_{0}]$ for some $a_{0}\in X_{+}$.
+Since $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{ob}(X)$ and $X_{+}$ is generating, $\bigcup\limits_{0\leq s\leq\varepsilon}{T_{s}}[-a_{0},a_{0}]\subseteq[-a_{1},a_{1}]$ for some $a_{1}\in X$.
+Similarly, for some $a_{2}\in X$, we have $\bigcup\limits_{0\leq s\leq\varepsilon}{T_{s}}[-a_{1},a_{1}]\subseteq[-a_{2},a_{2}]$ and
+$\bigcup\limits_{\varepsilon\leq s\leq 2\varepsilon}{T_{s}}[-a_{0},a_{0}]=T_{\varepsilon}\left(\bigcup\limits_{0\leq s\leq\varepsilon}{T_{s}}[-a_{0},a_{0}]\right)\subseteq T_{\varepsilon}[-a_{1},a_{1}]\subseteq[-a_{2},a_{2}].$
+It follows $\bigcup\limits_{0\leq s\leq 2\varepsilon}{T_{s}}[-a_{0},a_{0}]\subseteq[-(a_{1}+a_{2}),a_{1}+a_{2}]$. Repeating n-times the argument, we obtain $a_{1},a_{2},...,a_{n}\in X$ with
+$\bigcup\limits_{0\leq s\leq M}T_{s}[a,b]\subseteq\bigcup\limits_{0\leq s\leq n\varepsilon}{T_{s}}[-a_{0},a_{0}]\subseteq\left[-\sum_{i=1}^{n}a_{i},\sum_{i=1}^{n}a_{i}\right].$ (1)1( 1 )
+Since $[a,b]$ is arbitrary, (1) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{ob}(X)$, and since $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally order bounded.
+$ii)$ By Lemma 3.2, it follows from $i)$.
+$iii)$ Suppose $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{oc}(X)$ for some $\varepsilon>0$, and fix $M\geq 0$.
+Let $n\varepsilon\geq M$ and $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$.
+Pick a net $g^{(1)}_{{\beta_{(1)}}}\downarrow 0$ in $X$ such that, for each $\beta_{(1)}$ there is an $\alpha_{\beta_{(1)}}$ with ±$\pm T_{s}x_{\alpha}\leq g^{(1)}_{{\beta_{(1)}}}$ for $0\leq s\leq\varepsilon$ and $\alpha\geq\alpha_{\beta_{(1)}}$.
+Since $T_{\varepsilon}$ is order continuous, $T_{\varepsilon}x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ in $X$.
+It follows from $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{oc}(X)$ that there exists a net $g^{(2)}_{{\beta_{(2)}}}\downarrow 0$ in $X$ such that, for each $\beta_{(2)}$ there is $\alpha_{\beta_{(2)}}$ with ±$\pm T_{s}T_{\varepsilon}x_{\alpha}\leq g^{(2)}_{{\beta_{(2)}}}$ for $0\leq s\leq\varepsilon$ and $\alpha\geq\alpha_{\beta_{(2)}}$.
+Repeating n-times, we obtain nets
+$g^{(1)}_{{\beta_{(1)}}}\downarrow 0,\ g^{(2)}_{{\beta_{(2)}}}\downarrow 0,\ ...\ g^{(n)}_{{\beta_{(n)}}}\downarrow 0$
+so that, for each $\beta_{(i)}$, $i=1,2,...,n$, there is $\alpha_{\beta_{(i)}}$ satisfying
+±$\pm T_{s}T^{i-1}_{\varepsilon}x_{\alpha}\leq g^{(i)}_{{\beta_{(i)}}}\ \ \ \ (0\leq s\leq\varepsilon\ \ \text{\rm and}\ \ \alpha\geq\alpha_{\beta_{(i)}}).$
+Consider the net $g_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}=\sum_{i=1}^{n}g^{(i)}_{{\beta_{(i)}}}\downarrow 0$. Take an
+$\alpha_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}\geq\alpha_{\beta_{(1)}},\alpha_{\beta_{(2)}},....,\alpha_{\beta_{(n)}}.$
+Then, for each $\alpha\geq\alpha_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}$,
+±$\pm T_{s}T^{i-1}_{\varepsilon}x_{\alpha}\leq g^{(i)}_{{\beta_{(i)}}}\leq g_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}\ \ \ \ (0\leq s\leq\varepsilon\ \ \text{\rm and}\ \ i=1,2,...,n).$
+Therefore, for each $\alpha\geq\alpha_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}$,
+±$\pm T_{s}x_{\alpha}\leq g_{(\beta_{(1)},\beta_{(2)},...,\beta_{(n)})}\ \ \ \ (0\leq s\leq n\varepsilon).$ (2)2( 2 )
+Since $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ is arbitrary, (2) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq n\varepsilon}\in\text{\bf L}_{oc}(X)$, and hence $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{oc}(X)$.
+Since $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally order continuous.
+$iv)$ Suppose $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{onb}(X)$ for some $\varepsilon>0$ and fix $M\geq 0$.
+Let $n\varepsilon\geq M$ and take some $a\in X_{+}$.
+By Lemma 3.1, there exists $C\geq 0$ such that $\|T_{s}\|\leq C$ for all $0\leq s\leq\varepsilon$.
+Then
+$\sup\limits_{0\leq s\leq M}\|T_{s}a\|\leq\sup\limits_{0\leq s\leq n\varepsilon}\|T_{s}\|\|a\|\leq\Big{(}\sup\limits_{0\leq s\leq\varepsilon}\|T_{s}\|\Big{)}^{n}\|a\|\leq C\|a\|.$ (3)3( 3 )
+Since $a\in X_{+}$ is arbitrary, (3) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{onb}(X)$. Since $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally order to norm bounded. ∎
+In recent works [10, 11, 12] relatively uniformly continuous semigroups on VLs were deeply studied.
+The use of collectively qualified semigroups is to rough for getting progress in this direction.
+We postpone the investigation of relatively continuous semigroups on OVSs to a forthcoming paper.

### v3

@@
 Abstract
-It is proved that: each collectively order continuous set of operators between Archimedean ordered vector spaces with generating cones is collectively order bounded; and each collectively order to norm bounded set of operators from an ordered Banach space with a closed generating cone to a normed space is norm bounded.
+It is proved that: each collectively order continuous set of operators from an Archimedean OVS with a generating cone to an OVS is collectively order bounded; and each collectively order to norm bounded set of operators from an ordered Banach space with a closed generating cone to a normed space is norm bounded.
 Several applications to commutative operator semigroups on ordered vector spaces are given.
@@
 Recently, collectively qualified sets of operators on Banach latticed were investigate in [4, 8, 9, 10].
-The present paper generalizes results of [1, 7, 15] to the ordered vector space setting.
+The present paper generalizes previous results of [1, 6, 8, 16] to collectively qualified set of operators in the ordered vector space setting.
 The results of the paper are new even in the single operator case.
-For example, they imply that: 1) each order continuous operator between Archimedean ordered vector spaces is order bounded; 2) each order to norm continuous operator from an ordered Banach space with a closed generating cone to a normed space is continuous.
+For example, they imply that: 1) each order continuous operator on an Archimedean ordered vector space with a generating cone is order bounded; 2) each order to norm continuous operator on an ordered Banach space with a closed generating cone is bounded.
 We shall abbreviate a normed space (ordered vector space, ordered normed space, ordered Banach space, vector lattice, normed lattice, Banach lattice) by NS (OVS, ONS, OBS, VL, NL, BL).
-Throughout the text, vector spaces are real, operators are linear, symbol ${\cal L}(X,Y)$ (${\cal L}_{+}(X,Y)$, ${\cal L}_{ob}(X,Y)$, ${\cal L}_{oc}(X,Y)$, ${\cal L}_{onb}(X,Y)$) stands for the space of all (positive, order bounded, order continuous, order to norm bounded) operators from $X$ to $Y$.
-A symbol $x_{\alpha}\downarrow 0$ stands for a decreasing net in an OVS such that $\inf\limits_{\alpha}x_{\alpha}=0$.
+Throughout the text, vector spaces are real, operators are linear, a symbol ${\cal L}(X,Y)$ stands for the space of operators from $X$ to $Y$, and a symbol $x_{\alpha}\downarrow 0$ for a decreasing net in an OVS such that $\inf\limits_{\alpha}x_{\alpha}=0$.
 Quite often, working with collective properties of a set of operators in OVSs involves order convergent nets. Let $X$ be an OVS. A net $(x_{\alpha})$ in $X$ is
@@
     relative uniform convergent to $x\in X$ (briefly, ru-convergent to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{r}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices satisfying $-\frac{1}{n}u\leq x_{\alpha}-x\leq\frac{1}{n}u$ for all $\alpha\geq\alpha_{n}$.
-We shall use the following OVS version of [7, Definition 2.1].
+We shall use the following OVS version of [8, Definition 2.1].
 Let $X$ be an OVS, $B$ a set, $A$ a directed set, and =${\cal B}=\{(x^{b}_{\alpha})_{\alpha\in A}\}_{b\in B}$ a set of nets in $X$ indexed by $A$.
@@
 […]
-The following definition extends [7, Definition 2.1] to OVSs.
+The following definition extends [8, Definition 2.1] to OVSs.
 Definition 1.1.
@@
 The present paper is organized as follows.
-Section 2 begins with Theorem 2.1 saying that every collectively order continuous set of operators from an OVS to an Archimedean OVS is collectively order bounded.
+Section 2 begins with Theorem 2.1 saying that every collectively order continuous set of operators from an Archimedean OVS with a generating cone to an OVS is collectively order bounded.
 Then, we prove Theorem 2.2 that describes relationship between collectively r-continuous and collectively order bounded sets.
@@
 It follows, in particular that under these assumptions each operator which maps order interval into balls is bounded.
-Section 3 is devoted to applications of obtained results to commutative operator semigroups on OVSs.
+Section 3 is devoted to applications of obtained results to operator semigroups on OVSs.
 For the terminology and notations not explained in the text, we refer to [2, 3, 7, 14, 15].
 2 Main results
-We are going to prove the following results which generalize the results obtained recently in [7, 15] in the vector lattice setting to the OVS setting.
+We are going to prove the following results which generalize the results obtained in [6, 8, 16] in the vector lattice setting to the OVS setting.
   1. -
-    Every collectively order continuous set of operators from an Archimedean OVS with a generating cone to an OVS with a generating cone is collectively order bounded.
+    Every collectively order continuous set of operators from an Archimedean OVS with a generating cone to an OVS is collectively order bounded.
   1. -
@@
   1. -
-    Collectively r-continuous sets coincides with collectively order bounded sets when the OVS in the range has a generating cone.
+    Collectively r-continuous sets of operators between OVS with generating cones agree with collectively order bounded sets.
 2.1
-We begin with generalization of the first part of [7, Theorem 2.2]. The proof of the next theorem is based on ideas of the proof of the Abramovich–Sirotkin theorem in [1].
+We begin with generalization of the first part of [8, Theorem 2.2]. The proof of the next theorem is based on ideas of the proof of the Abramovich–Sirotkin theorem in [1].
 Theorem 2.1.
-Let $X$ and $Y$ be OVSs with generating cones and suppose $X$ is Archimedean. Then $\text{\bf L}_{oc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$.
+If $X$ is an Archimedean OVS with a generating cone. Then $\text{\bf L}_{oc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$ for each OVS $Y$.
 Proof.
 Let ${\cal T}\in\text{\bf L}_{oc}(X,Y)$.
-It is enough to prove order boundedness of ${\cal T}[0,b]$ for every $b\in X_{+}$.
+Since $X_{+}-X_{+}=X$, it is enough to prove order boundedness of ${\cal T}[0,b]$ for every $b\in X_{+}$.
 Take $[0,b]\subseteq X$.
@@
 Corollary 2.1.
-Each order continuous operator from an Archimedean OVS $X$ with generating cone to an OVS $Y$ with generating cone is order bounded.
-The following example shows that the Archimedean assumption, which has being missed in [1, Theorem 2.1], is essential in Corollary 2.1 even in the VL setting.
+Each order continuous operator from an Archimedean OVS $X$ with generating cone to an OVS $T$ is order bounded.
+The following example shows that the Archimedean assumption, which has being missed in [1, Theorem 2.1], is essential in Theorem 2.1 and Corollary 2.1 even in the VL setting.
 Example 2.1.
@@
 2.2
-Next, we generalize [15, Theorem 5.1] and [7, Theorem 2.2] to OVSs.
+Next, we generalize [6, Theorem 4.6.2], [16, Theorem 5.1], and [8, Theorem 2.2] to OVSs as follows.
 Theorem 2.2.
-Let $X$ and $Y$ be OVSs with generating cones then $\text{\bf L}_{rc}(X,Y)=\text{\bf L}_{ob}(X,Y)$.
+Let $X$ and $Y$ be OVSs.
+  1. $i)$
+    If $X$ has a generating cone then $\text{\bf L}_{rc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$.
+  2. $ii)$
+    If $Y$ has a generating cone then $\text{\bf L}_{ob}(X,Y)\subseteq\text{\bf L}_{rc}(X,Y)$.
+The idea of the proof follows the proof of [6, Theorem 4.6.2] with an additional use of the special net $x_{(k,z)}=\frac{1}{k}z$ form [1, Theorem 2.1] like in the proof of Theorem 2.1.
 Proof.
-Let ${\cal T}\in\text{\bf L}_{rc}(X,Y)$.
-In order to show ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ it suffices to prove that ${\cal T}[0,b]$ is order bounded in $Y$ for each $b\in X_{+}$.
-Take $[0,b]\subseteq X$.
+$i)$   Let ${\cal T}\in\text{\bf L}_{rc}(X,Y)$.
+Since $X_{+}-X_{+}=X$, for ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ it suffices to prove that ${\cal T}[0,b]$ is order bounded in $Y$ for each $b\in X_{+}$.
+Take $b\in X_{+}$.
 Consider =×${\cal I}=\mathbb{N}\times[0,b]$ and the net $x_{(k,z)}=\frac{1}{k}z$ like in the proof of Theorem 2.1.
@@
 It follows ${\cal T}\in\text{\bf L}_{ob}(X,Y)$.
-Now, suppose $Y=Y_{+}-Y_{+}$.
-Let ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ and $x_{\alpha}\operatorname{\xrightarrow[]{r}}0$.
+$ii)$   Let ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ and $x_{\alpha}\operatorname{\xrightarrow[]{r}}0$.
 Then, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices such that $-u\leq nx_{\alpha}\leq u$ (and hence $0\leq u+nx_{\alpha}\leq 2u$) for all $\alpha\geq\alpha_{n}$.
@@
 Indeed, the identity operator in the OBS $C^{1}[0,1]$ of continuously differentiable functions on $[0,1]$ is order bounded yet not order to norm bounded.
-The following result generalizes [7, Theorem 2.1] to OBSs.
+The following result generalizes [8, Theorem 2.1] to OBSs.
 Theorem 2.3.
@@
 More generally, any commutative operator semigroup $\mathcal{G}$ on $X$ can be written as $\mathcal{G}=(T_{S})_{S\in J}$, where $T_{S}=S$ and $J=\mathcal{G}$ is directed by a partial order $RS\succeq S$ for all $R,S\in\mathcal{G}$.
-The partial order $\succeq$ is extendable to the commutative operator semigroup ${\text{\rm co}}(\mathcal{G})$ via letting $RS\succeq S$ for all $R,S\in{\text{\rm co}}(\mathcal{G})$ (cf., [13, p.75]).
+The partial order $\succeq$ is extendable to the commutative operator semigroup ${\text{\rm co}}(\mathcal{G})$ via letting $RS\succeq S$ for all $R,S\in{\text{\rm co}}(\mathcal{G})$ (cf., [14, p.75]).
 Proposition 3.1.
@@
 In particular, $\{S\}_{S\succeq S_{0}}$ is collectively semicompact.
-Finally, we include the following convergence theorem.
+We include the following convergence theorem.
 Theorem 3.1.
@@
 Since $(Sx)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ lies in the weakly compact set $\overline{\text{\rm co}}(\bigcup_{S\in\mathcal{G}}S[-x,x])$, it has a weakly cluster point, say $y$.
-Then $\lim\limits_{S\to\infty}\|Sx-y\|=0$ by the Eberlein theorem (cf., [13, Theorem 1.5]).
+Then $\lim\limits_{S\to\infty}\|Sx-y\|=0$ by the Eberlein theorem (cf., [14, Theorem 1.5]).
 Since $x\in E_{+}$ is arbitrary, we conclude that $(S)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ converges strongly to a projection.
@@
 If $(T_{s})_{s\in J}$ is collectively almost order bounded then $(T_{s})_{s\in J}$ is mean ergodic.
-Uniformly continuous semigroups is a starting point in the theory of $C_{0}$-semigroups. Here we discuss similar elementary notions concerning semigroups on OVSs.
-The next two lemmas follow directly from [7, Theorems 2.1 and 2.2] and from Theorems 2.1, 2.2, and 2.3.
+Uniformly continuous semigroups is a starting point in the theory of $C_{0}$-semigroups.
+Here we discuss similar elementary notions concerning semigroups on OVSs.
+The next two lemmas follow directly from Theorems 2.1, 2.2, and 2.3.
 Lemma 3.1.
-Every collectively order to norm bounded set of operators on an (OBS) with a closed generating cone is norm bounded.
+Every collectively order to norm bounded set of operators on an OBS with a closed generating cone is norm bounded.
 Lemma 3.2.
@@
 In recent works [11, 12, 13] relatively uniformly continuous semigroups on VLs were deeply studied.
-The use of collectively qualified semigroups is to rough for getting progress in this direction.
+The use of collectively qualified semigroups is too rough for getting progress in this direction.
 We postpone the investigation of relatively continuous semigroups on OVSs to a forthcoming paper.

### v4

@@
 Abstract
-It is proved that: each collectively order continuous set of operators from an Archimedean OVS with a generating cone to an OVS is collectively order bounded; and each collectively order to norm bounded set of operators from an ordered Banach space with a closed generating cone to a normed space is norm bounded.
+It is proved that: each collectively order continuous set of operators from an Archimedean ordered vector space with a generating cone to an ordered vector space is collectively order bounded; and each collectively order to norm bounded set of operators from an ordered Banach space with a closed generating cone to a normed space is norm bounded.
 Several applications to commutative operator semigroups on ordered vector spaces are given.
+Eduard Emelyanov1, Nazife Erkurşun-Özcan2, Svetlana Gorokhova 3
+$1$ Sobolev Institute of Mathematics, Novosibirsk, Russia
+$2$ Hacettepe University, Ankara, Turkey
+$3$ Southern Mathematical Institute, Vladikavkaz, Russia
 Keywords: Ordered vector space, collectively qualified set of operators, commutative operator semigroup
@@
 1 Introduction
-Collectively compact sets of operators between normed spaces were introduced by P.
-M.
-Anselone and T.
-W.
-Palmer [5].
-Recently, collectively qualified sets of operators on Banach latticed were investigate in [4, 8, 9, 10].
-The present paper generalizes previous results of [1, 6, 8, 16] to collectively qualified set of operators in the ordered vector space setting.
-The results of the paper are new even in the single operator case.
-For example, they imply that: 1) each order continuous operator on an Archimedean ordered vector space with a generating cone is order bounded; 2) each order to norm continuous operator on an ordered Banach space with a closed generating cone is bounded.
-We shall abbreviate a normed space (ordered vector space, ordered normed space, ordered Banach space, vector lattice, normed lattice, Banach lattice) by NS (OVS, ONS, OBS, VL, NL, BL).
-Throughout the text, vector spaces are real, operators are linear, a symbol ${\cal L}(X,Y)$ stands for the space of operators from $X$ to $Y$, and a symbol $x_{\alpha}\downarrow 0$ for a decreasing net in an OVS such that $\inf\limits_{\alpha}x_{\alpha}=0$.
-Quite often, working with collective properties of a set of operators in OVSs involves order convergent nets. Let $X$ be an OVS. A net $(x_{\alpha})$ in $X$ is
+Collectively compact sets of operators between normed spaces were introduced by Anselone and Palmer [5].
+Recently, collectively qualified sets of operators were investigated in [4, 10, 11, 12, 14].
+In the present paper, we generalize some results of [1, 8, 10, 18] to collectively qualified sets in the ordered vector space setting.
+Results of the paper are new even in the single operator case.
+For example, they imply that: 1) each order continuous operator on an Archimedean ordered vector space with a generating cone is order bounded; 2) each order to norm bounded operator on an ordered Banach space with a closed generating cone is continuous.
+Throughout the text, vector spaces are real, operators are linear, ${\cal L}(X,Y)$ stands for the space of operators from a vector space $X$ to a vector space $Y$, and $x_{\alpha}\downarrow 0$ stands for a decreasing net in an ordered vector space such that $\inf\limits_{\alpha}x_{\alpha}=0$.
+Using of order convergence is common in study of collective properties in ordered vector spaces. A net $(x_{\alpha})$ in an ordered vector space $X$ is
   1. -
-    order convergent to $x\in X$ (briefly, o-convergent to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{o}}x$) whenever there exists a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$ there is an $\alpha_{\beta}$ satisfying $-g_{\beta}\leq x_{\alpha}-x\leq g_{\beta}$ for all $\alpha\geq\alpha_{\beta}$.
+    order convergent to $x\in X$ (briefly, o-convergent to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{o}}x$) whenever there exists a net $g_{\beta}\downarrow 0$ in $X$ such that, for each $\beta$ there is $\alpha_{\beta}$ satisfying $-g_{\beta}\leq x_{\alpha}-x\leq g_{\beta}$ for all $\alpha\geq\alpha_{\beta}$.
   2. -
-    relative uniform convergent to $x\in X$ (briefly, ru-convergent to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{r}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices satisfying $-\frac{1}{n}u\leq x_{\alpha}-x\leq\frac{1}{n}u$ for all $\alpha\geq\alpha_{n}$.
-We shall use the following OVS version of [8, Definition 2.1].
-Let $X$ be an OVS, $B$ a set, $A$ a directed set, and =${\cal B}=\{(x^{b}_{\alpha})_{\alpha\in A}\}_{b\in B}$ a set of nets in $X$ indexed by $A$.
+    relative uniform convergent to $x\in X$ (briefly, ru-convergent to $x$, or $x_{\alpha}\operatorname{\xrightarrow[]{ru}}x$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices satisfying $-\frac{1}{n}u\leq x_{\alpha}-x\leq\frac{1}{n}u$ for all $\alpha\geq\alpha_{n}$.
+We shall use the following version of [10, Definition 1.2].
+Let $X$ be an ordered vector space, $B$ a set, $A$ a directed set, and =${\cal B}=\{(x^{b}_{\alpha})_{\alpha\in A}\}_{b\in B}$ a set of nets in $X$ indexed by $A$.
 We say that ${\cal B}$
   1. -
-    collective o-converges to 00 (briefly, ${\cal B}\operatorname{\xrightarrow[]{c-o}}0$) if there exists a net $g_{\beta}\downarrow 0$ such that, for each $\beta$, there is an $\alpha_{\beta}$ satisfying ±$\pm x^{b}_{\alpha}\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$ and $b\in B$.
+    collective o-converges to 00 (briefly, ${\cal B}\operatorname{\xrightarrow[]{c-o}}0$) if there exists a net $g_{\beta}\downarrow 0$ such that, for each $\beta$, there is $\alpha_{\beta}$ satisfying ±$\pm x^{b}_{\alpha}\leq g_{\beta}$ for $\alpha\geq\alpha_{\beta}$ and $b\in B$.
   2. -
-    collective ru-converges to 00 (briefly, ${\cal B}\operatorname{\xrightarrow[]{c-r}}0$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices with ±$\pm x^{b}_{\alpha}\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$ and $b\in B$.
+    collective ru-converges to 00 (briefly, ${\cal B}\operatorname{\xrightarrow[]{c-ru}}0$) if, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices with ±$\pm x^{b}_{\alpha}\leq\frac{1}{n}u$ for $\alpha\geq\alpha_{n}$ and $b\in B$.
 […]
-The following definition extends [8, Definition 2.1] to OVSs.
 Definition 1.1.
-Let ${\cal T}\subseteq{\cal L}(X,Y)$.
-When $X$ and $Y$ are OVSs,
+(cf., [10, Definition 2.1]) Let ${\cal T}\subseteq{\cal L}(X,Y)$.
+When $X$ and $Y$ are ordered vector spaces, the set ${\cal T}$ is
+  1. $a)$
+    collectively order bounded (briefly, ${\cal T}\in{\bf L}_{ob}(X,Y)$) if, for every order interval $[a,b]$ in $X$ there exists an order interval $[u,v]$ in $Y$ such that ${\cal T}[a,b]:=\bigcup\limits_{T\in\mathcal{T}}{T}[a,b]\subseteq[u,v]$.
+  2. $b)$
+    collectively o-continuous (briefly, ${\cal T}\in\text{\bf L}_{oc}(X,Y)$) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c-o}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$;
+  3. $c)$
+    collectively ru-continuous (briefly, ${\cal T}\in\text{\bf L}_{ruc}(X,Y)$) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c-ru}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$;
+When $X$ is an ordered vector space and $Y$ an ordered normed space, the set ${\cal T}$ is
+  1. $d)$
+    collectively almost order bounded (briefly, ${\cal T}\in\text{\bf L}_{aob}(X,Y)$) if, for every $[a,b]\subseteq X$ and $\varepsilon>0$, there exists $[u,v]\subseteq Y$ such that ${\cal T}[a,b]\subseteq[u,v]+\varepsilon B_{Y}$.
+When $X$ is an ordered vector space and $Y$ is a normed space, the set ${\cal T}$ is
+  1. $e)$
+    collectively order to norm bounded (briefly, ${\cal T}\in\text{\bf L}_{onb}(X,Y)$) if the set ${\cal T}[a,b]$ is norm bounded for every $[a,b]\subseteq X$.
+The present paper is organized as follows.
+Section 2 begins with Theorem 2.1 saying that every collectively order continuous set of operators from an Archimedean ordered vector space with a generating cone to an ordered vector space is collectively order bounded.
+Theorem 2.4 describes relationship between collectively ru-continuous and collectively order bounded sets, and Theorem 2.8 says that, under rather mild assumptions, every collectively order to norm bounded set of operators is norm bounded.
+As a consequence, Corollary 2.9 says that each order to norm bounded operator from an ordered Banach space with a closed generating cone to a normed space is continuous.
+Proposition 2.10 tell us that every order bounded operator between ordered Banach spaces with closed generating cones is continuous.
+Section 3 is devoted to applications of obtained results to operator semigroups on ordered vector spaces.
+For the terminology and notations that are not explained in the text, we refer to [2, 3, 9, 17].
+2 Main results
+We are going to prove the following results which generalize the corresponding ones obtained in [1, 10, 18] to the setting of ordered vector spaces.
   1. -
-    ${\cal T}$ is collectively order bounded (briefly, ${\cal T}\in{\bf L}_{ob}(X,Y)$) if, for every order interval $[a,b]$ in $X$ there exists an order interval $[u,v]$ in $Y$ such that ${\cal T}[a,b]:=\bigcup\limits_{T\in\mathcal{T}}{T}[a,b]\subseteq[u,v]$.
+    Theorem 2.1 saying that every collectively order continuous set of operators from an Archimedean ordered vector space with a generating cone to any ordered vector space is collectively order bounded.
   1. -
-    ${\cal T}$ is collectively o-continuous (briefly, ${\cal T}\in\text{\bf L}_{oc}(X,Y)$) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c-o}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$;
+    Theorem 2.4 saying that collectively ru-continuous sets of operators between ordered vector spaces with generating cones agree with collectively order bounded sets.
   1. -
-    ${\cal T}$ is collectively ru-continuous (briefly, ${\cal T}\in\text{\bf L}_{rc}(X,Y)$) if ${\cal T}x_{\alpha}\operatorname{\xrightarrow[]{c-r}}0$ whenever $x_{\alpha}\operatorname{\xrightarrow[]{r}}0$;
-When $X$ is an OVS and $Y$ is an ONS,
-  1. -
-    ${\cal T}$ is collectively almost order bounded (briefly, ${\cal T}\in\text{\bf L}_{aob}(X,Y)$) if, for every $[a,b]\subseteq X$ and $\varepsilon>0$, there exists $[u,v]\subseteq Y$ such that ${\cal T}[a,b]\subseteq[u,v]+\varepsilon B_{Y}$.
-When $X$ is a normed space and $Y$ is an ONS,
-  1. -
-    ${\cal T}$ is collectively semicompact (briefly, ${\cal T}\in\text{\bf L}_{sc}(X,Y)$) if, for every $\varepsilon>0$, there exists $[u,v]\subseteq Y$ such that ${\cal T}B_{X}\subseteq[u,v]+\varepsilon B_{Y}$.
-When $X$ is an OVS and $Y$ is a normed space,
-  1. -
-    ${\cal T}$ is collectively order to norm bounded (briefly, ${\cal T}\in\text{\bf L}_{onb}(X,Y)$) if the set ${\cal T}[a,b]$ is norm bounded for every $[a,b]\subseteq X$.
-The present paper is organized as follows.
-Section 2 begins with Theorem 2.1 saying that every collectively order continuous set of operators from an Archimedean OVS with a generating cone to an OVS is collectively order bounded.
-Then, we prove Theorem 2.2 that describes relationship between collectively r-continuous and collectively order bounded sets.
-Finally, we prove Theorem 2.3 saying that under rather mild assumptions, every collectively order to norm bounded set of operators is norm bounded.
-It follows, in particular that under these assumptions each operator which maps order interval into balls is bounded.
-Section 3 is devoted to applications of obtained results to operator semigroups on OVSs.
-For the terminology and notations not explained in the text, we refer to [2, 3, 7, 14, 15].
-2 Main results
-We are going to prove the following results which generalize the results obtained in [6, 8, 16] in the vector lattice setting to the OVS setting.
-  1. -
-    Every collectively order continuous set of operators from an Archimedean OVS with a generating cone to an OVS is collectively order bounded.
-  1. -
-    Every collectively order to norm bounded set of operators from an OBS with a closed generating cone to a NS is norm bounded
-  1. -
-    Collectively r-continuous sets of operators between OVS with generating cones agree with collectively order bounded sets.
-2.1
-We begin with generalization of the first part of [8, Theorem 2.2]. The proof of the next theorem is based on ideas of the proof of the Abramovich–Sirotkin theorem in [1].
+    Theorem 2.8 saying that every collectively order to norm bounded set of operators from an ordered Banach space with a closed generating cone to a normed space norm bounded
+Let us begin with the following generalization of the first part of [10, Theorem 2.2] which is a collective version of the result by Abramovich and Sirotkin.
+The proof is based on ideas in the proof of [1, Theorem 2.1].
 Theorem 2.1.
-If $X$ is an Archimedean OVS with a generating cone. Then $\text{\bf L}_{oc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$ for each OVS $Y$.
+Let $X$ be an Archimedean ordered vector space with a generating cone. Then $\text{\bf L}_{oc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$ for every ordered vector space $Y$.
 Proof.
@@
 Therefore, ${\cal T}[0,b]\subseteq[-(m_{\beta_{0}}+1)g_{\beta_{0}},(m_{\beta_{0}}+1)g_{\beta_{0}}]$. ∎
-In particular, Theorem 2.1 generalizes the Abramovich–Sirotkin theorem [1, Theorem 2.1] as follows.
+In particular, this extends the Abramovich–Sirotkin theorem as follows.
 Corollary 2.2.
-Each order continuous operator from an Archimedean OVS $X$ with generating cone to an OVS $T$ is order bounded.
-The following example shows that the Archimedean assumption, which has being missed in [1, Theorem 2.1], is essential in Theorem 2.1 and Corollary 2.1 even in the VL setting.
+Each order continuous operator from an Archimedean ordered vector space with generating cone to any ordered vector space is order bounded.
+Note that the Archimedean assumption, which has being missed in [1, Theorem 2.1], is essential in Theorem 2.1 and in Corollary 2.2 even in the vector lattice setting.
 Example 2.3.
-Let $X$ be the lexicographically ordered real plane $\mathbb{R}^{2}$, $Y$ the pointwise ordered $\mathbb{R}^{2}$, and $T:X\to Y$ the identity operator.
-Both $X$ and $Y$ are VLs.
-The operator $T$ is not order bounded since the order interval $[(0,0),(2,0)]$ of $X$ contains line $x=1$ that is not order bounded in $Y$.
+Let $X$ be the lexicographically ordered real plane $\mathbb{R}^{2}$, $Y$ the coordinate-wise ordered $\mathbb{R}^{2}$.
+Both $X$ and $Y$ are vector lattices.
+The identity operator $T:X\to Y$ is not order bounded since the order interval $[(0,0),(2,0)]$ of $X$ contains line $x=1$ that is not order bounded in $Y$.
 However, $T$ is order continuous.
@@
 It follows that $T(x_{\alpha},y_{\alpha})\operatorname{\xrightarrow[]{o}}0$ in $Y$.
-2.2
-Next, we generalize [6, Theorem 4.6.2], [16, Theorem 5.1], and [8, Theorem 2.2] to OVSs as follows.
+An inspection of the proof of [1, Theorem 2.1] shows that Abramovich and Sirotkin proved already that each ru-continuous operator between vector lattices $E$ and $F$ is order bounded.
+Conversely, each order bounded operator $T:E\to F$ is ru-continuous since, for every $u\in E_{+}$ there exists $w\in F$ such that $T[-u,u]\subseteq[-w,w]$, and hence $|x_{\alpha}|\leq\frac{1}{n}u$ implies $|Tx_{\alpha}|\leq\frac{1}{n}w$.
+This assertion was independently proved by Taylor and Troitsky in [18, Theorem 5.1].
+Note that this fact can also be derived from the nonstandard criteria of order boundedness of operators [8, Theorem 1.7.2] obtained by the first author.
+The following theorem generalizes this fact to the setting of ordered vector space.
 Theorem 2.4.
-Let $X$ and $Y$ be OVSs.
+Let $X$ and $Y$ be ordered vector spaces.
   1. $i)$
-    If $X$ has a generating cone then $\text{\bf L}_{rc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$.
+    If $X$ has a generating cone then $\text{\bf L}_{ruc}(X,Y)\subseteq\text{\bf L}_{ob}(X,Y)$.
   2. $ii)$
-    If $Y$ has a generating cone then $\text{\bf L}_{ob}(X,Y)\subseteq\text{\bf L}_{rc}(X,Y)$.
-The idea of the proof follows the proof of [6, Theorem 4.6.2] with an additional use of the special net $x_{(k,z)}=\frac{1}{k}z$ form [1, Theorem 2.1] like in the proof of Theorem 2.1.
+    If $Y$ has a generating cone then $\text{\bf L}_{ob}(X,Y)\subseteq\text{\bf L}_{ruc}(X,Y)$.
 Proof.
-$i)$   Let ${\cal T}\in\text{\bf L}_{rc}(X,Y)$.
-Since $X_{+}-X_{+}=X$, for ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ it suffices to prove that ${\cal T}[0,b]$ is order bounded in $Y$ for each $b\in X_{+}$.
+$i)$   Let ${\cal T}\in\text{\bf L}_{ruc}(X,Y)$.
+Since $X_{+}-X_{+}=X$, to show ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ it suffices to prove that ${\cal T}[0,b]$ is order bounded in $Y$ for each $b\in X_{+}$.
 Take $b\in X_{+}$.
@@
 If $(m,z)\geq(n,b)$ in ${\cal I}$ then either $(m,z)=(n,b)$ or $m>n$, and hence $0\leq x_{(m,z)}=\frac{1}{m}z\leq\frac{1}{n}b$.
-Thus, $x_{(m,z)}\operatorname{\xrightarrow[]{r}}0$ in $X$.
-Since ${\cal T}\in\text{\bf L}_{rc}(X,Y)$, there exist $u\in Y_{+}$ and an increasing sequence $(k_{n},z_{n})$ in ${\cal I}$ satisfying ±$\pm T\big{(}\frac{z}{k}\big{)}=\pm Tx_{(k,z)}\leq\frac{1}{n}u$ for all $T\in{\cal T}$ and $(k,z)\geq(k_{n},z_{n})$.
+Thus, $x_{(m,z)}\operatorname{\xrightarrow[]{ru}}0$ in $X$.
+Since ${\cal T}\in\text{\bf L}_{ruc}(X,Y)$, there exist $u\in Y_{+}$ and an increasing sequence $(k_{n},z_{n})$ in ${\cal I}$ satisfying ±$\pm T\big{(}\frac{z}{k}\big{)}=\pm Tx_{(k,z)}\leq\frac{1}{n}u$ for all $T\in{\cal T}$ and $(k,z)\geq(k_{n},z_{n})$.
 In particular, ±$\pm T\big{(}\frac{z}{k_{1}+1}\big{)}=\pm Tx_{(k_{1}+1,z)}\leq u$ for all $T\in{\cal T}$ and $z\in[0,b]$.
@@
 It follows ${\cal T}\in\text{\bf L}_{ob}(X,Y)$.
-$ii)$   Let ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ and $x_{\alpha}\operatorname{\xrightarrow[]{r}}0$.
-Then, for some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices such that $-u\leq nx_{\alpha}\leq u$ (and hence $0\leq u+nx_{\alpha}\leq 2u$) for all $\alpha\geq\alpha_{n}$.
-Since ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ and $Y=Y_{+}-Y_{+}$, there exist $w_{1},w_{2}\in Y_{+}$ satisfying $Tu\in[-w_{1},w_{1}]$ and $T(u+nx_{\alpha})\in[-w_{2},w_{2}]$ for all $T\in{\cal T}$ and $\alpha\geq\alpha_{n}$.
-Then $nTx_{\alpha}\in[-(w_{1}+w_{2}),(w_{1}+w_{2})]$ (and hence ±$\pm Tx_{\alpha}\leq\frac{w_{1}+w_{2}}{n}$) for $T\in{\cal T}$ and $\alpha\geq\alpha_{n}$.
-It follows ${\cal T}\in\text{\bf L}_{rc}(X,Y)$.
+$ii)$   Let ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ and $x_{\alpha}\operatorname{\xrightarrow[]{ru}}0$.
+For some $u\in X_{+}$, there exists an increasing sequence $(\alpha_{n})$ of indices such that $nx_{\alpha}\in[-u,u]$ for all $\alpha\geq\alpha_{n}$.
+Since ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ and $Y=Y_{+}-Y_{+}$ then ${\cal T}[-u,u]\subseteq[-w,w]$ for some $w\in Y_{+}$, and hence $n{\cal T}x_{\alpha}\subseteq[-w,w]$ for all $\alpha\geq\alpha_{n}$.
+It follows ${\cal T}\in\text{\bf L}_{ruc}(X,Y)$.
 ∎
 Corollary 2.5.
-Order bounded operators between OVSs with generating cones agree with r-continuous operators.
-2.3
-Recall that order intervals in an OBS $Y$ with a closed cone are norm bounded iff the cone is normal (cf., [3, Theorem 2.40]). We include the following elementary collective extension of this fact .
+Order bounded operators between ordered vector spaces with generating cones agree with ru-continuous operators.
+Recall that order intervals in an ordered Banach space with a closed cone are norm bounded if and only if the cone is normal (cf., [3, Theorem 2.40]).
+We include the following elementary collective extension of this fact.
 Proposition 2.6.
-Let $X$ be an OVS and $Y$ be an OBS with closed generating normal cone. Then $\text{\bf L}_{aob}(X,Y)\subseteq\text{\bf L}_{onb}(X,Y)$.
+Let $X$ be an ordered vector space and $Y$ an ordered Banach space with closed generating normal cone. Then $\text{\bf L}_{aob}(X,Y)\subseteq\text{\bf L}_{onb}(X,Y)$.
 Proof.
-Let ${\cal T}\in\text{\bf L}_{aob}(X,Y)$ and take $[a,b]\subseteq X$.
-Then $\bigcup_{T\in\mathcal{T}}{T}[a,b]\subseteq[-y_{1},y_{1}]+B_{Y}$ for some $y_{1}\in Y_{+}$.
-Since $Y_{+}$ is normal, $[-y_{1},y_{1}]+B_{Y}$ is norm bounded by [3, Theorem 2.40].
+Let ${\cal T}\in\text{\bf L}_{aob}(X,Y)$ and $[a,b]\subseteq X$.
+Then ${\cal T}[a,b]\subseteq[-y,y]+B_{Y}$ for some $y\in Y_{+}$.
+Since $Y_{+}$ is normal, $[-y,y]+B_{Y}$ is norm bounded by [3, Theorem 2.40].
 Therefore, ${\cal T}\in\text{\bf L}_{onb}(X,Y)$.
 ∎
-The normality of $Y_{+}$ is essential in Proposition 2.1.
-Indeed, the identity operator in the OBS $C^{1}[0,1]$ of continuously differentiable functions on $[0,1]$ is order bounded yet not order to norm bounded.
-The following result generalizes [8, Theorem 2.1] to OBSs.
+The normality of positive cone in the range is essential in Proposition 2.6.
+Example 2.7.
+Let $X=C^{1}[0,1]$ be an ordered Banach space of continuously differentiable functions on $[0,1]$.
+The identity operator $I_{X}$ is norm and order bounded yet not order to norm bounded.
+In particular, $I_{X}\in\text{\bf L}_{aob}(X,Y)\setminus\text{\bf L}_{onb}(X,Y)$.
+The following result generalizes [10, Theorem 2.1] to ordered Banach spaces.
 Theorem 2.8.
-Let $X$ be an OBS with a closed generating cone and $Y$ be a normed space. Then each ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ is norm bounded.
+Let $X$ be an ordered Banach space with a closed generating cone and $Y$ be a normed space. Then each ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ is norm bounded.
 Proof.
 Let ${\cal T}\in\text{\bf L}_{onb}(X,Y)$.
-Suppose in contrary ${\cal T}B_{X}$ is not norm bounded.
-The M.
-G.
-Krein – Smulian theorem (cf., [3, Theorem 2.37]) implies $\alpha B_{X}\subseteq B_{X}\cap X_{+}-B_{X}\cap X_{+}$ for some $\alpha>0$, and hence ${\cal T}(B_{X}\cap X_{+})$ is not norm bounded.
+Suppose, in contrary, ${\cal T}B_{X}$ is not norm bounded.
+The Krein – Smulian theorem (cf., [3, Theorem 2.37]) implies $\alpha B_{X}\subseteq B_{X}\cap X_{+}-B_{X}\cap X_{+}$ for some $\alpha>0$, and hence ${\cal T}(B_{X}\cap X_{+})$ is not norm bounded.
 Pick sequences $(x_{n})$ in $B_{X}\cap X_{+}$ and $(T_{n})$ in ${\cal T}$ with $\|T_{n}x_{n}\|\geq n^{3}$ for all $n$, and set $x:=\|\cdot\|$-$\sum\limits_{n=1}^{\infty}n^{-2}x_{n}\in X_{+}$.
@@
 ∎
-It is worth noting that even in the case of a single operator, Theorem 2.3 generalizes the well known fact that every order bounded operator between BLs is bounded.
-The condition $X_{+}-X_{+}=X$ is essential in Theorem 2.3.
+Corollary 2.9.
+Every order to norm bounded operator from an ordered Banach space $X$ with a closed generating cone to a normed space $Y$ is continuous.
+It is worth noting that Corollary 2.9 generalizes the well known fact that every order bounded operator between Banach lattices is bounded.
+It is also known that every positive operator $T:X\to Y$ is continuous whenever $X$ is an ordered Banach space with a closed generating cone and $Y$ is an ordered Banach space with a closed cone, see Arendt [6, Appendix], Aliprantis and Tourky [3, Theorem 2.32], or Arendt and Nittka [7, Theorem 2.8].
+Note that a positive operator need not to be order to norm bounded.
+Indeed, the identity operator $I_{X}$ on an ordered Banach space $X$ is order to norm bounded if and only if the positive cone $X_{+}$ is normal.
+Therefore, the next proposition does not follow from Corollary 2.9.
+Its proof below is a modification of the proof of [3, Theorem 2.32].
+Proposition 2.10.
+Let $X$ and $Y$ be ordered Banach spaces with closed generating cones. Then every order bounded operator $T:X\to Y$ is continuous.
+Proof.
+Let $T:X\to Y$ be order bounded.
+To establish that $T$ is continuous, it suffices to show that $T$ has a closed graph.
+So, assume $x_{n}\operatorname{\xrightarrow[]{\|\cdot\|}}0$ in $X$ and $Tx_{n}\operatorname{\xrightarrow[]{\|\cdot\|}}y$ in $Y$.
+We must show that $y=0$.
+According to [3, Lemma 2.30], there exist a subsequence $(n_{k})$ such that $x_{n_{k}}\operatorname{\xrightarrow[]{ru}}0$.
+By Corollary 2.5, $T$ is ru-continuous, and hence $Tx_{n_{k}}\operatorname{\xrightarrow[]{ru}}0$.
+Thus, there exist some $u\in Y$ and a further subsequence $(n_{k_{m}})$ such that ±$\pm Tx_{n_{k_{m}}}\leq\frac{1}{m}u$, and hence $\frac{1}{m}u\pm Tx_{n_{k_{m}}}\in Y_{+}$ for each $m$.
+Since $Y_{+}$ is closed and $\frac{1}{m}u\pm Tx_{n_{k_{m}}}\operatorname{\xrightarrow[]{\|\cdot\|}}\pm y$, we conclude that ±$\pm y\in Y_{+}$.
+That is, y = 0 and the proof is finished.
+∎
+We have the following collective extension of Proposition 2.10.
+Proposition 2.11.
+Let $X$ and $Y$ be ordered Banach spaces with closed generating cones, and $X^{\sim}\neq\{0\}$. The following assertions are equivalent.
+  1. $i)$
+    Every ${\cal T}\in\text{\bf L}_{ob}(X,Y)$ is norm bounded.
+  2. $ii)$
+    The cone $Y_{+}$ is normal.
+Proof.
+$i)\Longrightarrow ii)$ Assume that $Y_{+}$ is not normal.
+Then, there exists $y_{0}\in X_{+}$ such that the interval $[0,y_{0}]$ is not bounded.
+Pick $f_{0}\neq 0$ in $X^{\sim}$ and take $x_{0}\in X$ with $f_{0}(x_{0})=1$.
+Then, $\{f_{0}\otimes y\}_{y\in[0,y_{0}]}\in\text{\bf L}_{ob}(X,Y)$.
+However, the set $\{f_{0}\otimes y\}_{y\in[0,y_{0}]}$ is not norm bounded because of $\bigcup\limits_{y\in[0,y_{0}]}(f_{0}\otimes y)(x_{0})=[0,y_{0}]$.
+$ii)\Longrightarrow i)$ Let ${\cal T}\in\text{\bf L}_{ob}(X,Y)$.
+Then, ${\cal T}\in\text{\bf L}_{aob}(X,Y)$.
+So, ${\cal T}\in\text{\bf L}_{onb}(X,Y)$ by Proposition 2.6.
+It follows from Theorem 2.8 that ${\cal T}$ is norm bounded.
+∎
+The following example shows that the condition that $X_{+}$ is closed and generating is essential in Theorem 2.8, in Corollary 2.9, and in Proposition 2.10.
 Example 2.12.
-Take any infinite dimensional Banach space $X$ with the trivial cone $\{0\}$, consider arbitrary discontinuous operator $T$ on $X$, and observe that $T$ is order to norm bounded simply because order intervals in $X$ are singletons.
+Consider an infinite dimensional Banach space $X$.
+  1. $a)$
+    Let $X_{+}:=\{0\}$ and take an arbitrary discontinuous operator $T$ on $X$.
+Then $T:(X,X_{+})\to X$ is order to norm bounded and order bounded simply because order intervals in $X$ are singletons.
+Clearly $X_{+}$ is closed but not generating.
+  2. $b)$
+    Pick a normalized Hamel basis $\{\xi:\xi\in\Xi\}$ of $X$, identify $X$ with $c_{00}(\Xi)$, and let $X_{+}:=\Big{\{}\sum\limits_{\xi\in\Xi}\lambda_{\xi}\xi:\lambda_{\xi}\geq 0\Big{\}}$.
+Then $X_{+}-X_{+}=X$ but $X_{+}$ is not closed ((((e.g., by [7, Proposition 4.1])))).
+Pick any sequence $(\xi_{n})_{n=0}^{\infty}$ of distinct elements of $\Xi$ and define $T:(X,X_{+})\to X$ by $T\xi=0$ for $\xi\in\Xi\setminus\{\xi_{n}:n\in\mathbb{N}\}$ and $T\xi_{n}=n\xi_{0}$ for $n\in\mathbb{N}$.
+Clearly, $T$ is order to norm bounded and positive, yet not continuous.
+Example 2.13.
+Let $X$ be an ordered vector space with a generating cone which satisfies the dominated decomposition property ((((i.e., $[0,v_{1}]+[0,v_{2}]=[0,v_{1}+v_{2}]$ for every $v_{1},v_{2}$ in $X_{+}$)))), and let $Y$ be a Dedekind complete vector lattice.
+It is well known that the ordered vector space ${\cal L}_{ob}(X,Y)$ of order bounded operators of ${\cal L}(X,Y)$ is a Dedekind complete vector lattice (cf., [19, Theorem 83.4]).
+Let $A$ be an order bounded subset of ${\cal L}_{ob}(X,Y)$, say $A\subseteq[-T,T]$ for some positive operator $T\in{\cal L}_{ob}(X,Y)$.
+Let $x\in X_{+}$.
+Since $|Sz|\leq|S|z\leq Tz\leq Tx$ for all $z\in[0,x]$ and $S\in A$, then the set $A$ is collectively order bounded.
+From the other hand, not every collectively order bounded set of operators is order bounded.
+Indeed, consider $X=\ell^{\infty}$, $Y=\mathbb{R}$, and $A=\{\delta^{n}:n\in\mathbb{N}\}\subset\ell^{\infty}_{+}$, where $\delta^{n}(\mathbb{x})=x_{n}$.
+The set $A$ is collectively order bounded, yet there is no $f\in\ell^{\infty}_{+}$ with $A\subseteq[0,f]$, since otherwise $f(1,1,\ldots,1,\ldots)\geq\sum_{k=1}^{n}\delta^{k}(1,1,\ldots,1,\ldots)=n$ for every $n\in\mathbb{N}$ that is absurd.
 3 Some applications to operator semigroups
-We shall use the notation $(T_{s})_{s\in J}$ for a one-parameter operator semigroup on a vector space $X$ (with $J=\mathbb{R}_{+}$ in the continuous-parameter case and $J=\mathbb{Z}_{+}$ in the discrete-parameter case).
+Under a semigroup $(T_{s})_{s\geq 0}$ on a vector space $X$ we understand a family of operators on $X$ satisfying $T_{s+t}=T_{s}T_{t}$ for all $s,t\in\mathbb{R}_{+}$, and $T_{0}=I_{X}$, where $I_{X}$ is the identity operator on $X$.
+We shall also use the notation $(T_{s})_{s\in J}$ for a one-parameter operator semigroup on $X$, where $J=\mathbb{R}_{+},\mathbb{R},\mathbb{Z}_{+}$, or $\mathbb{Z}$.
+Such a semigroup is always commutative.
 More generally, any commutative operator semigroup $\mathcal{G}$ on $X$ can be written as $\mathcal{G}=(T_{S})_{S\in J}$, where $T_{S}=S$ and $J=\mathcal{G}$ is directed by a partial order $RS\succeq S$ for all $R,S\in\mathcal{G}$.
-The partial order $\succeq$ is extendable to the commutative operator semigroup ${\text{\rm co}}(\mathcal{G})$ via letting $RS\succeq S$ for all $R,S\in{\text{\rm co}}(\mathcal{G})$ (cf., [14, p.75]).
+The partial order $\succeq$ is extendable to the commutative operator semigroup ${\text{\rm co}}(\mathcal{G})$ via letting $RS\succeq S$ for all $R,S\in{\text{\rm co}}(\mathcal{G})$ (cf., [17, p.75]).
 Proposition 3.1.
-Let $\mathcal{G}$ be a collectively order to norm bounded commutative operator semigroup on an OBS $X$ with a closed generating cone containing a (weakly) compact operator $S_{0}$.
+Let $\mathcal{G}$ be a collectively order to norm bounded commutative operator semigroup on an ordered Banach space $X$ with a closed generating cone containing a (weakly) compact operator $S_{0}$.
 Then the semigroup $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively (weakly) compact.
@@
 Since $S_{0}B_{X}$ is relatively (weakly) compact, $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively (weakly) compact. ∎
-Combining Propositions 3.1 and 2.1, and keeping in mind that every OBS with a generating normal cone is Archimedean, we obtain.
-Corollary 3.1.
-Let $\mathcal{G}$ be a collectively almost order bounded commutative operator semigroup on an OBS $X$ with a closed, generating, and normal cone.
-If $\mathcal{G}$ contains a (weakly) compact operator $S_{0}$ then $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively (weakly) compact.
-In particular, $\{S\}_{S\succeq S_{0}}$ is collectively (weakly) compact.
-Next, we consider commutative operator semigroups on AL-spaces.
-Proposition 3.2.
-Let $\mathcal{G}$ be a collectively almost order bounded commutative operator semigroup on on an AL-space $E$ containing a weakly compact operator $S_{0}$.
-Then the semigroup $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively semicompact.
-In particular, $\{S\}_{S\succeq S_{0}}$ is collectively semicompact.
-Proof.
-By Theorem 2.3, $\|S\|\leq M<\infty$ for $S\in\mathcal{G}$.
-Then $\|S\|\leq M<\infty$ for all $S\in{\text{\rm co}}(\mathcal{G})$.
-We conclude $\bigcup_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}SB_{E}\subseteq MS_{0}B_{E}$.
-Since the set $S_{0}B_{E}$ is relatively weakly compact, it is almost order bounded by the Dunford-Pettis theorem (cf., [15][Theorem 2.5.4]).
-Thus, $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively semicompact.
-∎
-Combining Propositions 3.2 and 2.1 we obtain.
-Corollary 3.2.
-Let $\mathcal{G}$ be a collectively almost order bounded commutative operator semigroup on an AL-space $E$ containing a weakly compact operator $S_{0}$.
-Then the semigroup $(S)_{{\text{\rm co}}(\mathcal{G})\ni S\succeq S_{0}}$ is collectively semicompact.
-In particular, $\{S\}_{S\succeq S_{0}}$ is collectively semicompact.
-We include the following convergence theorem.
 Theorem 3.2.
-Let $\mathcal{G}$ be a collectively almost order bounded commutative operator semigroup on an order continuous BL $E$.
+Let $\mathcal{G}$ be a collectively almost order bounded commutative operator semigroup on an ordered Banach space $X$ with a closed generating cone, and assume that every order interval in $X$ is weakly compact.
 Then the net $(S)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ converges strongly to a projection.
 Proof.
-By Theorem 2.3, $\|S\|\leq M<\infty$ for $S\in\mathcal{G}$, and hence for all $S\in\overline{{\text{\rm co}}}(\mathcal{G})$.
-Let $x\in E_{+}$.
-By the assumption, the set $\bigcup_{S\in\mathcal{G}}S[-x,x]$ is almost order bounded.
-Since $E$ is order continuous, $\bigcup_{S\in\mathcal{G}}S[-x,x]$ is relatively weakly compact due to the Grothendieck lemma (cf., [2, Theorem 3.44]).
-Since $(Sx)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ lies in the weakly compact set $\overline{\text{\rm co}}(\bigcup_{S\in\mathcal{G}}S[-x,x])$, it has a weakly cluster point, say $y$.
-Then $\lim\limits_{S\to\infty}\|Sx-y\|=0$ by the Eberlein theorem (cf., [14, Theorem 1.5]).
-Since $x\in E_{+}$ is arbitrary, we conclude that $(S)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ converges strongly to a projection.
+Since order intervals in $X$ are weakly compact then every almost order bounded subset of $X$ is relatively weakly compact due to the Grothendieck result (cf., [2, Theorem 3.44]).
+In particular, $\mathcal{G}$ be a collectively order to norm bounded.
+By Theorem 2.8, $\|S\|\leq M$ for some $M<\infty$ and all $S\in\mathcal{G}$.
+Let $x\in X_{+}$.
+Then, the set $\bigcup_{S\in\mathcal{G}}S[-x,x]$ is almost order bounded, and hence relatively weakly compact.
+Since $(Sx)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ lies in the weakly compact set $\overline{\text{\rm co}}(\bigcup_{S\in\mathcal{G}}S[-x,x])$, it has a weak cluster point, say $y$.
+Then $\lim\limits_{S\to\infty}\|Sx-y\|=0$ by the Eberlein theorem (cf., [17, Theorem 1.5]).
+Since $x\in X_{+}$ is arbitrary and $X_{+}$ is generating, $(S)_{S\in\overline{{\text{\rm co}}}(\mathcal{G})}$ converges strongly to a projection.
 ∎
 Corollary 3.3.
-Let $(T_{s})_{s\in J}$ be either a $C_{0}$-semigroup or a one-parameter discrete operator semigroup on an order continuous BL $E$.
+Let $(T_{s})_{s\in J}$ be either a $C_{0}$-semigroup or one-parameter discrete operator semigroup on an order continuous Banach lattice $E$.
 If $(T_{s})_{s\in J}$ is collectively almost order bounded then $(T_{s})_{s\in J}$ is mean ergodic.
-Uniformly continuous semigroups is a starting point in the theory of $C_{0}$-semigroups.
-Here we discuss similar elementary notions concerning semigroups on OVSs.
-The next two lemmas follow directly from Theorems 2.1, 2.2, and 2.3.
-Lemma 3.1.
-Every collectively order to norm bounded set of operators on an OBS with a closed generating cone is norm bounded.
-Lemma 3.2.
-Let $X$ be and OVS with generating cone. Then $\text{\bf L}_{rc}(X)=\text{\bf L}_{ob}(X)$. If $X$ is additionally Archimedean then $\text{\bf L}_{oc}(X)\subseteq\text{\bf L}_{ob}(X)$.
-Next definition localizes several notions concerning semigrops.
+The following example shows that the order continuity of the norm in $E$ is essential in Corollary 3.3.
+Example 3.4.
+(cf., [9, Exercise 2.2.22]) Consider the operator $T:c\to c$ such that $T(a_{1},a_{2},\ldots a_{k},\ldots)=\big{(}\frac{1}{2}a_{1},\frac{2}{3}a_{2},\ldots,\frac{k}{k+1}a_{k},\ldots\big{)}$.
+Clearly, $\big{\{}T^{n}\big{\}}_{n=1}^{\infty}\in\text{\bf L}_{ob}(c)$ for some $\varepsilon>0$.
+However, the Cesàro averages $\frac{1}{n}\sum_{j=0}^{n-1}T^{j}$ do not converge in the norm at $(1,1,\ldots 1,\ldots)\in c$.
+The following corollary of Theorem 3.2 is immediate, because order intervals in the predual of a Neumann algebra are weakly compact by the Akemann theorem (cf., [9, Theorem 3.3.3]).
+Corollary 3.5.
+Let $(T_{s})_{s\in J}$ be either a $C_{0}$-semigroup or one-parameter discrete operator semigroup on the predual $\mathcal{M}_{*}$ of a von Neumann algebra $\mathcal{M}$.
+If $(T_{s})_{s\in J}$ is collectively almost order bounded then $(T_{s})_{s\in J}$ is mean ergodic.
+Uniformly continuous semigroups is a starting point in the theory of $C_{0}$-semigroups. Here we discuss similar elementary notions concerning semigroups on ordered vector spaces.
 Definition 3.6.
-A semigroup $(T_{s})_{s\geq 0}$ on an OVS $X$ is
+A semigroup $(T_{s})_{s\geq 0}$ on an ordered vector space $X$ is
   1. $a)$
-    locally order bounded if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{ob}(X)$ for every $M$.
+    locally collectively order bounded ((((shortly, lco-bounded))))
+    whenever $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{ob}(X)$ for every $M>0$.
   2. $b)$
-    locally o-continuous if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{oc}(X)$ for every $M$.
+    lco-continuous whenever $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{oc}(X)$ for every $M>0$.
   3. $c)$
-    locally ru-continuous if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{rc}(X)$ for every $M$.
-A semigroup $(T_{s})_{s\geq 0}$ on a normed OVS $X$ is
-  1. $e)$
-    locally order to norm bounded if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{onb}(X)$
-    for every $M$.
-In several important cases, in each of the four items of Definition 3.1 the condition “for every $M$” can be replaced by “for some $M$”.
+    lcru-continuous whenever $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{ruc}(X)$ for every $M>0$.
+A semigroup $(T_{s})_{s\geq 0}$ on a normed ordered vector space $X$ is
+  1. $d)$
+    lcon-bounded if $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{onb}(X)$ for every $M>0$.
+In several important cases, the condition “for every $M>0$” can be replaced by “for some $\varepsilon>0$” in each of the four items of Definition 3.6.
 Proposition 3.7.
-Let $(T_{s})_{s\geq 0}$ be a semigroup on an OVS $X$ such that $X_{+}-X_{+}=X$.
+Let ${\cal T}=(T_{s})_{s\geq 0}$ be a semigroup on an ordered vector space $X$ such that $X_{+}-X_{+}=X$.
   1. $i)$
-    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{ob}(X)$ for some $\varepsilon>0$, then $(T_{s})_{s\geq 0}$ is locally order bounded.
+    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{ob}(X)$ for some $\varepsilon>0$, then ${\cal T}$ is lco-bounded.
   2. $ii)$
-    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{rc}(X)$ for some $\varepsilon>0$, then $(T_{s})_{s\geq 0}$ is locally ru-continuous.
+    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{ruc}(X)$ for some $\varepsilon>0$, then ${\cal T}$ is lcru-continuous.
   3. $iii)$
-    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{oc}(X)$ for some $\varepsilon>0$, then $(T_{s})_{s\geq 0}$ is locally order continuous.
+    If $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{oc}(X)$ for some $\varepsilon>0$, then ${\cal T}$ is lco-continuous.
   4. $iv)$
-    If $X$ is Banach space, $X_{+}$ is closed, and $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{onb}(X)$ for some $\varepsilon>0$, then $(T_{s})_{s\geq 0}$ is locally order to norm bounded.
+    If $X$ is Banach space, $X_{+}$ is closed, and $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{onb}(X)$ for some $\varepsilon>0$, then ${\cal T}$ is lcon-bounded.
 Proof.
@@
 $\bigcup\limits_{0\leq s\leq M}T_{s}[a,b]\subseteq\bigcup\limits_{0\leq s\leq n\varepsilon}{T_{s}}[-a_{0},a_{0}]\subseteq\left[-\sum_{i=1}^{n}a_{i},\sum_{i=1}^{n}a_{i}\right].$ (1)1( 1 )
-Since $[a,b]$ is arbitrary, (1) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{ob}(X)$, and since $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally order bounded.
-$ii)$ By Lemma 3.2, it follows from $i)$.
+Since $[a,b]$ is arbitrary, (1) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{ob}(X)$, and since $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally collectively order bounded.
+$ii)$ By Theorem 2.4, it follows from $i)$.
 $iii)$ Suppose $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{oc}(X)$ for some $\varepsilon>0$, and fix $M\geq 0$.
@@
 Since $x_{\alpha}\operatorname{\xrightarrow[]{o}}0$ is arbitrary, (2) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq n\varepsilon}\in\text{\bf L}_{oc}(X)$, and hence $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{oc}(X)$.
-Since $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally order continuous.
+Since $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally collectively order continuous.
 $iv)$ Suppose $\big{\{}T_{s}\big{\}}_{0\leq s\leq\varepsilon}\in\text{\bf L}_{onb}(X)$ for some $\varepsilon>0$ and fix $M\geq 0$.
-Let $n\varepsilon\geq M$ and take some $a\in X_{+}$.
-By Lemma 3.1, there exists $C\geq 0$ such that $\|T_{s}\|\leq C$ for all $0\leq s\leq\varepsilon$.
+Take $n\in\mathbb{N}$ with $n\varepsilon\geq M$.
+Let $a\in X_{+}$.
+By Theorem 2.1, there exists $C\geq 0$ such that $\|T_{s}\|\leq C$ for all $0\leq s\leq\varepsilon$.
 Then
-$\sup\limits_{0\leq s\leq M}\|T_{s}a\|\leq\sup\limits_{0\leq s\leq n\varepsilon}\|T_{s}\|\|a\|\leq\Big{(}\sup\limits_{0\leq s\leq\varepsilon}\|T_{s}\|\Big{)}^{n}\|a\|\leq C\|a\|.$ (3)3( 3 )
-Since $a\in X_{+}$ is arbitrary, (3) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{onb}(X)$. Since $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally order to norm bounded. ∎
-In recent works [11, 12, 13] relatively uniformly continuous semigroups on VLs were deeply studied.
-The use of collectively qualified semigroups is too rough for getting progress in this direction.
-We postpone the investigation of relatively continuous semigroups on OVSs to a forthcoming paper.
+$\sup\limits_{0\leq s\leq M}\|T_{s}a\|\leq\sup\limits_{0\leq s\leq n\varepsilon}\|T_{s}\|\|a\|\leq\Big{(}\sup\limits_{0\leq s\leq\varepsilon}\|T_{s}\|\Big{)}^{n}\|a\|\leq C^{n}\|a\|.$ (3)3( 3 )
+Since $a\in X_{+}$ is arbitrary, (3) implies $\big{\{}T_{s}\big{\}}_{0\leq s\leq M}\in\text{\bf L}_{onb}(X)$.
+As $M\geq 0$ is arbitrary, $(T_{s})_{s\geq 0}$ is locally collectively order to norm bounded.
+∎
+Under rather mild assumptions, every order bounded (o-continuous, ru-continuous, order to norm bounded) operator between ordered Banach spaces gives a rise to lco-bounded (resp., lco-bounded, lco-continuous, lcru-continuous, lcon-bounded) operator semigroup $(T_{t})_{t\geq 0}$.
+Example 3.8.
+Let $S:X\to Y$ be a bounded operator between Banach spaces.
+Let $Z=X\oplus Y$, define a bounded operator $A:Z\to Z$ by $A(x,y)=(0,Sx)$, and set $T_{t}=e^{tA}=\|\cdot\|$-$\sum_{k=0}^{\infty}\frac{t^{k}A^{k}}{k!}$ for $t\geq 0$.
+Then $(T_{t})_{t\geq 0}$ is a uniformly continuous semigroup on the Banach space $Z$ such that $T_{t}=I+tA$ for $t\geq 0$ because $A^{2}=0$.
+Suppose $X$ and $Y$ are ordered Banach spaces and $X_{+}-X_{+}=X$.
+  1. $(a)$
+    Let $Y_{+}$ be normal and $S:X\to Y$ be order bounded.
+Then $S$ is order to norm bounded since $S$ is order bounded and $Y_{+}$ is normal.
+Theorem 2.8 implies that $S$, and hence $A$ is bounded.
+Since $A$ is order bounded, it follows from $T_{t}=I+tA$ for $t\geq 0$ that $(T_{t})_{t\geq 0}$ is lco-bounded.
+  2. $(b)$
+    Let $Y_{+}$ be normal and $S:X\to Y$ be ru-continuous.
+By Theorem 2.4, $S$ is order bounded, and hence $S$ is order to norm bounded because $Y_{+}$ is normal.
+Due to Theorem 2.8, $S$ is bounded.
+Thus $A$ is bounded.
+Since $A$ is ru-continuous, $T_{t}=I+tA$ implies $(T_{t})_{t\geq 0}$ is lcru-continuous.
+  3. $(c)$
+    Suppose additionally that $X$ is Archimedean.
+Let $Y_{+}$ be normal and $S:X\to Y$ be order continuous.
+By Theorem 2.1, $S$ is order bounded.
+Since $Y_{+}$ is normal, $S$ is order to norm bounded.
+Theorem 2.8 implies boundedness of $S$, and hence of $A$.
+Since $A$ is order continuous and $T_{t}=I+tA$, we have $(T_{t})_{t\geq 0}$ is lco-continuous.
+  4. $(d)$
+    Let $S:X\to Y$ be order to norm bounded. Then, by Theorem 2.8, $S$ is bounded, and so is $A$. Since $A$ is order to norm bounded, $T_{t}=I+tA$ implies $(T_{t})_{t\geq 0}$ is lcon-bounded.
+In recent works of Glück, Kandić, Kaplin, and Kramar Fijavž [13, 15, 16] relatively uniformly continuous semigroups on vector lattices were deeply studied.
+Using of collectively qualified semigroups is too rough for getting progress in this direction.
+We postpone the investigation of relatively continuous semigroups on ordered vector spaces to a forthcoming paper.